### v1

# Equivariant Machine Learning on Graphs with Nonlinear Spectral Filters

Abstract
Equivariant machine learning is an approach for designing deep learning models that respect the symmetries of the problem, with the aim of reducing model complexity and improving generalization. In this paper, we focus on an extension of shift equivariance, which is the basis of convolution networks on images, to general graphs. Unlike images, graphs do not have a natural notion of domain translation. Therefore, we consider the graph functional shifts as the symmetry group: the unitary operators that commute with the graph shift operator. Notably, such symmetries operate in the signal space rather than directly in the spatial space. We remark that each linear filter layer of a standard spectral graph neural network (GNN) commutes with graph functional shifts, but the activation function breaks this symmetry. Instead, we propose nonlinear spectral filters (NLSFs) that are fully equivariant to graph functional shifts and show that they have universal approximation properties. The proposed NLSFs are based on a new form of spectral domain that is transferable between graphs. We demonstrate the superior performance of NLSFs over existing spectral GNNs in node and graph classification benchmarks.
1 Introduction
In many fields, such as chemistry [Gilmer et al., 2017], biology [Gaudelet et al., 2021, Barabasi and Oltvai, 2004], social science [Borgatti et al., 2009], and computer graphics [Wang et al., 2019], data can be described by graphs. In recent years, there has been a tremendous interest in the development of machine learning models for graph-structured data [Wu et al., 2020, Bronstein et al., 2017]. This young field, often called graph machine learning (graph-ML) or graph representation learning, has made significant contributions to the applied sciences, e.g., in protein folding, molecular design, and drug discovery [Jumper et al., 2021, Atz et al., 2021, Méndez-Lucio et al., 2021, Stokes et al., 2020], and has impacted the industry with applications in social media, recommendation systems, traffic prediction, computer graphics, and natural language processing, among others.
Geometric Deep Learning (GDL) [Bronstein et al., 2017] is a design philosophy for machine learning models where the model is constructed to inherently respect symmetries present in the data, aiming to reduce model complexity and enhance generalization. By incorporating knowledge of these symmetries into the model, it avoids the need to expend parameters and data to learn them. This inherent respect for symmetries is automatically generalized to test data, thereby improving generalization [Bietti et al., 2021, Petrache and Trivedi, 2024, Elesedy and Zaidi, 2021]. For instance, convolutional neural networks (CNNs) respect the translation symmetries of 2D images, with weight sharing due to these symmetries contributing significantly to their success [Krizhevsky et al., 2012]. Respecting node re-indexing in a scalable manner revolutionized machine learning on graphs [Bronstein et al., 2017, 2021] and has placed GNNs as a main general-purpose tool for processing graph-structured data. Moreover, within GNNs, respecting the 3D Euclidean symmetries of the laws of physics (rotations, reflections, and translations) led to state-of-the-art performance in molecule processing [Duval et al., 2024].
Our Contribution. In this paper, we focus on GDL for graph-ML. We consider extensions of shift symmetries from images to general graphs. Since graphs do not have a natural notion of domain translation, as opposed to images, we propose considering functional translations instead. We model the group of translations on graphs as the group of all unitary operators on signals that commute with the graph shift operator. Such unitary operators are called graph functional shifts. Note that each linear filter layer of a standard spectral GNN commutes with graph functional shifts, but the activation function breaks this symmetry. Instead, we propose non-linear spectral filters (NLSFs) that are fully equivariant to graph functional shifts and have universal approximation properties.
In Sec. 3, we introduce our NLSFs based on new notions of analysis and synthesis that map signals between their node-space representations and spectral representations. Our transforms are related to standard graph Fourier and inverse Fourier transforms but differ from them in one important aspect. One key property of our analysis transform is that it is independent of a specific choice of Laplacian eigenvectors. Hence, our spectral representations are transferable between graphs. In comparison, standard graph Fourier transforms are based on an arbitrary choice of eigenvectors, and therefore, the standard frequency domain is not transferable. To achieve transferability, standard graph Fourier methods resort to linear filter operations based on functional calculus. Since we do not have this limitation, we can operate on the frequency coefficients with arbitrary nonlinear functions such as multilayer perceptrons. In Sec. 4, we present theoretical results of our NLSFs, including the universal approximation and expressivity properties. In Sec. 5, we demonstrate the efficacy of our NLSFs in node and graph classification benchmarks, where our method outperforms existing spectral GNNs.
2 Background
Notation. For $N\in\mathbb{N}$, we denote $[N]=\{1,\ldots,N\}$. We denote matrices by boldface uppercase letter $\mathbf{B}$, vectors (assumed to be columns) by lowercase boldface $\mathbf{b}$, and the entries of matrices and vectors are denoted with the same letter in lower case, e.g., $\mathbf{B}=(b_{i,j})_{i,j\in[N]}$. Let $G=([N],\mathcal{E},\mathbf{A},\mathbf{X})$ be an undirected graph with a node set $[N]$, an edge set $\mathcal{E}\subset[N]\times[N]$, an adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ representing the edge weights, and a node feature matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ containing $d$-dimensional node attributes. Let $\mathbf{D}$ be the diagonal degree matrix of $G$, where the diagonal element $D_{i,i}$ is the degree of node $i$. Denote by $\bm{\Delta}$ any normal graph shift operator (GSO). For example, $\bm{\Delta}$ could be the combinatorial graph Laplacian or the normalized graph Laplacian given by $\mathbf{L}=\mathbf{D}-\mathbf{A}$ and $\mathbf{N}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}$, respectively. Let $\mathbf{\Delta}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{\top}$ be the eigendecomposition of $\mathbf{\Delta}$, where $\mathbf{V}$ is the eigenvector matrix and $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{N})$ is the diagonal matrix with eigenvalues $(\lambda_{i})_{i=1}^{N}$ ordered by $\left\lvert\lambda_{1}\right\rvert\leq\ldots\leq\left\lvert\lambda_{N}\right\rvert$. An eigenspace is the span of all eigenvectors corresponding to the same eigenvalue. Let $\mathbf{P}_{i}=\mathbf{P}_{\bm{\Delta};i}$ denote the projection upon the $i$-th eigenspace of ${\bm{\Delta}}$ in increasing order of $\left\lvert\lambda\right\rvert$. We denote the Euclidean norm by $\left\lVert{\mathbf{X}}\right\rVert_{2}$. We define the channel-wise signal norm $\left\lVert{\mathbf{X}}\right\rVert_{\rm sig}$ of a matrix ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ as the vector $\left\lVert{\mathbf{X}}\right\rVert_{\rm sig}=(\left\lVert{\mathbf{X}}_{:,j}\right\rVert_{2})_{j=1}^{d}\in{\mathbb{R}}^{d}$, where ${\mathbf{X}}_{:,j}$ is the $j$-th column on ${\mathbf{X}}$. We abbreviate multilayer perceptrons by MLP.
2.1 Linear Graph Signal Processing
Spectral GNNs define convolution operators on graphs via the spectral domain. Given a self-adjoint graph shift operator (GSO) $\bm{\Delta}$, e.g., a graph Laplacian, the Fourier modes of the graph are defined to be the eigenvectors $\{\mathbf{v}_{i}\}_{i=1}^{N}$ of $\bm{\Delta}$ and the eigenvalues $\{\lambda_{i}\}_{i=1}^{N}$ are the frequencies. A spectral filter is defined to directly satisfy the “convolution theorem” [Bracewell and Kahn, 1966] for graphs. Namely, given a signal ${\mathbf{X}}\in{\mathbb{R}}^{N\times d}$ and a function $\mathbf{Q}:{\mathbb{R}}\rightarrow{\mathbb{R}}^{d^{\prime}\times d}$, the operator $\mathbf{Q}(\bm{\Delta}):\mathbb{R}^{N\times d}\rightarrow\mathbb{R}^{N\times d^{\prime}}$ defined by
$\mathbf{Q}(\bm{\Delta}){\mathbf{X}}:=\sum_{i=1}^{N}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}{\mathbf{X}}\mathbf{Q}(\lambda_{i})^{\top},$ (1)
is called a filter. Here, $d^{\prime}$ is the number of output channels. Spectral GNNs, e.g., [Defferrard et al., 2016b, Kipf and Welling, 2016, Levie et al., 2018, Bianchi et al., 2021], are graph convolutional networks where convolutions are via Eq. (1), with a trainable function $\mathbf{Q}$ at each layer, and a nonlinear activation function.
2.2 Equivariant GNNs
Equivariance describes the ability of functions $f$ to respect symmetries. It is expressed as $f(H_{\kappa}x)=H_{\kappa}f(x)$, where $\mathcal{K}\ni\kappa\mapsto H_{\kappa}$ is an action of a symmetry group $\mathcal{K}$ on the domain of $f$. GNNs [Scarselli et al., 2008, Bruna et al., 2013], including spectral GNNs [Defferrard et al., 2016a, Kipf and Welling, 2016] and subgraph GNNs [Frasca et al., 2022, Bevilacqua et al., 2022], are inherently permutation equivariant w.r.t. the ordering of nodes. This means that the network’s operations are unaffected by the specific arrangement of nodes, a property stemming from passive symmetries [Villar et al., 2023] where transformations are applied to both the graph signals and the graph domain. This permutation equivariance is often compared to the translation equivariance in CNNs [LeCun et al., 1995, 2010], which involves active symmetries [Huang et al., 2024]. The key difference between the two symmetries lies in the domain: while CNNs operate on a fixed domain with signals transforming within it, graphs lack a natural notion of domain translation. To address this, we consider graph functional shifts as the symmetry group, defined by unitary operators that commute with the graph shift operator. This perspective allows for our NLSFs to be interpreted as an extension of active symmetry within the graph context, bridging the gap between the passive and active symmetries inherent to GNNs and CNNs, respectively.
3 Nonlinear Spectral Graph Filters
In this section, we present new concepts of analysis and synthesis under which the spectral domain is transferrable between graphs. Following these concepts, we introduce new GNNs that are equivariant to functional symmetries – symmetries of the Hilbert space of signals rather than symmetries in the domain of definition of the signal [Lim et al., 2024].
3.1 Translation Equivariance of CNNs and GNNs
For motivation, we start with the grid graph $R$ with node set $[M]^{2}$ and circular adjacency $\mathbf{B}$, we define the translation operator $\mathbf{T}_{m,n}$ by $[m,n]$ as
$\mathbf{T}_{m,n}\in{\mathbb{R}}^{M^{2}\times M^{2}};\quad(\mathbf{T}_{m,n}\mathbf{x})_{i,j}=\mathbf{x}_{l,k}\quad\text{where}\quad l=(i-m)\ {\rm mod\ }M\ \ \ \ \text{and}\ \ \ k=(j-n)\ {\rm mod\ }M\ .$
Note that any $\mathbf{T}_{m,n}$ is a unitary operator that commutes with the grid Laplacian $\bm{\Delta}_{R}$, i.e., $\mathbf{T}_{m,n}\bm{\Delta}_{R}=\bm{\Delta}_{R}\mathbf{T}_{m,n}$, and therefore it belongs to the group of all unitary operators $\mathcal{U}_{R}$ that commute with the grid Laplacian $\bm{\Delta}_{R}$. In fact, the space of isotropic convolution operators (with 90o rotation and reflection symmetric filters) can be seen as the space of all normal operators [^1] The operator $\mathbf{B}$ is normal iff $\mathbf{B}^{*}\mathbf{B}=\mathbf{B}\mathbf{B}^{*}$. Equivalently, iff $\mathbf{B}$ has an orthogonal eigendecomposition. that commute with unitary operators from $\mathcal{U}_{R}$ [Cohen and Welling, 2016]. Applying a non-linearity after the convolution retains this equivariance, and hence, we can build multi-layer CNNs that commute with $\mathcal{U}_{R}$. By the universal approximation theorem [Cybenko, 1989, Funahashi, 1989, Leshno et al., 1993], this allows us to approximate any continuous function that commutes with $\mathcal{U}_{R}$.
Note that such translation equivariance cannot be extended to general graphs. Achieving equivariance to graph functional shifts through linear spectral convolutional layers $\mathbf{Q}(\bm{\Delta})$ is straightforward, since these layers commute with the space of all unitary operators $\mathcal{U}_{{\bm{\Delta}}}$ that commute with $\bm{\Delta}$. However, introducing non-linearity $\rho(\mathbf{Q}(\bm{\Delta})\mathbf{X})$ breaks the symmetry. That is, there exists $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$ such that
$\rho\left(\mathbf{Q}(\bm{\Delta})\mathbf{U}\mathbf{X}\right)\neq\mathbf{U}\rho\left(\mathbf{Q}(\bm{\Delta})\mathbf{X}\right).$
This means that multi-layer spectral GNNs do not commute with $\mathcal{U}_{G}$, and are hence not appropriate as approximators of general continuous functions that commute with $\mathcal{U}_{G}$. Instead, we propose in this paper a multi-layer GNN that is fully equivariant to $\mathcal{U}_{G}$, which we show to be universal: it can approximate any continuous graph-signal function (w.r.t. some metric) commuting with $\mathcal{U}_{G}$.
3.2 Graph Functional Symmetries and Their Relaxations
We define the symmetry group of graph functional shifts as follows.
Definition 3.1 (Graph Functional Shifts).
The space of graph functional shifts is the unitary subgroup $\mathcal{U}_{{\bm{\Delta}}}$, where a unitary matrix $\mathbf{U}$ is in $\mathcal{U}_{{\bm{\Delta}}}$ iff it commutes with the GSO ${\bm{\Delta}}$, namely, $\mathbf{U}\bm{\Delta}=\bm{\Delta}\mathbf{U}$.
It is important to note that functional shifts, in general, are not induced from node permutations. Instead, functional shifts are related to the notion of functional maps [Ovsjanikov et al., 2012] used in shape correspondence.
A fundamental challenge with the symmetry group in Def. 3.1 is its lack of transferability between different graphs. Hence, we propose to relax this symmetry group. Let $g_{1},\ldots,g_{S}:\mathbb{R}\rightarrow\mathbb{R}$ be the indicator functions of the intervals $\{[l_{s},l_{s+1}]\}_{s=1}^{S}$, which constitute a partition of the frequency band $[l_{1},l_{S}]\subset\mathbb{R}$. The operators $g_{j}(\bm{\Delta})$, interpreted via functional calculus Eq. (1), are projections of the signal space upon band-limited signals. Namely, $g_{j}(\bm{\Delta})=\sum_{i:\lambda_{i}\in[l_{j},l_{j+1}]}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}$. In our work, we consider filters $g_{j}$ that are supported on the dyadic sub-bands $\left[\lambda_{N}r^{S-j+1},\lambda_{N}r^{S-j}\right]$, where $10<r<$ 1 is the decay rate. See Fig. 3 in App. C for an illustrated example. Note that for $j=1$, the sub-band falls in $[0,\lambda_{N}r^{S-1}]$. The total band $[0,l_{S}]$ is $[0,\lambda_{N}]$.
Definition 3.2 (Relaxed Functional Shifts).
The space of relaxed functional shifts with respect to the filter bank $\{g_{j}\}_{j=1}^{K}$ (of indicators) is the unitary subgroup $\mathcal{U}^{g}_{{\bm{\Delta}}}$, where a unitary matrix $\mathbf{U}$ is in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ iff it commutes with $g_{j}({\bm{\Delta}})$ for all $j$, namely, $\mathbf{U}g_{j}(\bm{\Delta})=g_{j}(\bm{\Delta})\mathbf{U}$.
Similarly, we can relax functional shifts by restricting to the leading eigenspaces.
Definition 3.3 (Leading Functional Shifts).
The space of leading functional shifts is the unitary subgroup $\mathcal{U}^{J}_{{\bm{\Delta}}}$, where a unitary $\mathbf{U}$ is in $\mathcal{U}^{J}_{{\bm{\Delta}}}$ iff it commutes with the eigenspace projections $\{\mathbf{P}_{j}\}_{j=1}^{J}$.
3.3 Analysis and Synthesis
We use the terminology of analysis and synthesis, as in signal processing [Mallat, 1989], to describe transformations of signals between their graph and spectral representations. Here, we consider two settings: the eigenspace projections case and the filter bank (of indicators) case.
Spectral Index Case. We first define the analysis and synthesis using the spectral index by
$\mathcal{A}^{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\Big{(}\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}\Big{)}_{j=1}^{J}\in{\mathbb{R}}^{J\times d}\text{\ \ \ and \ \ \ }\mathcal{S}^{\text{ind}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\sum_{j=1}^{J}\mathbf{r}_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (2)
respectively, where in synthesis, the product and division are element-wise. Here, $\mathbf{R}=(\mathbf{r}_{j})_{j=1}^{J}\in{\mathbb{R}}^{J\times d}$ are the spectral coefficients to be synthesized, and $0\leq a\leq 1$ and $0<e\ll 1$ are parameters that promote stability. The term index stems from the fact that eigenvalues are treated according to their index when defining the projections $\mathbf{P}_{j}$. Note that the synthesis here differs from classic signal processing as it depends on a given signal on the graph. When treating ${\bm{\Delta}}$ and ${\mathbf{X}}$ as fixed, this synthesis operation is denoted by $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{\text{ind}}(\mathbf{R}):=\mathcal{S}^{\text{ind}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})$. We similarly denote $\mathcal{A}^{\text{ind}}_{{\bm{\Delta}}}({\mathbf{X}}):=\mathcal{A}^{\text{ind}}({\bm{\Delta}},{\mathbf{X}})$.
Filter Bank Case. We also consider and define the analysis and synthesis in the filter bank by
$\mathcal{A}^{\text{val}}({\bm{\Delta}},{\mathbf{X}})=\Big{(}\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\Big{)}_{j=1}^{K}\in{\mathbb{R}}^{K\times d}\text{\ and \ }\mathcal{S}^{\text{val}}(\mathbf{R};{\bm{\Delta}},{\mathbf{X}})=\sum_{j=1}^{K}\mathbf{r}_{j}\ \frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (3)
respectively, where in synthesis, the product and division are element-wise. Here, $\mathbf{R}=(\mathbf{r}_{j})_{j=1}^{K}\in{\mathbb{R}}^{K\times d}$ are the spectral coefficients to be synthesized, and $a,e$ are as before. In App. A.3, we show that the synthesis operator is stably invertible. The term value refers to how eigenvalues are used based on their magnitude when defining the projections $g_{j}({\bm{\Delta}})$. As before, we denote $\mathcal{S}^{\text{val}}_{{\bm{\Delta}},{\mathbf{X}}}$ and $\mathcal{A}^{\text{val}}_{{\bm{\Delta}}}$.
3.4 Definitions of Nonlinear Spectral Filters
Figure 1: Illustration of nonlinear spectral filters for equivariant machine learning on graphs. Given a graph $G$, the node features $\mathbf{X}$ are projected onto eigenspaces (analysis $\mathcal{A}$). The function $\Psi$ map a sequence of frequency responses to a sequence of frequency responses. The signal is represented in the graph domain using the nonlinear spectral coefficients (synthesis $\mathcal{S}$).
We introduce three novel types of non-linear spectral filters (NLSF): Node-level NLSFs, Graph-level NLSFs, and Pooling-NLSFs. Fig. 1 illustrates our NLSFs for equivariant machine learning on graphs.
Node-level NLSFs. To be able to transfer NLSFs between different graphs and signals, one key property of NLSF is that they do not depend on the specific basis chosen in each eigenspace. This independence is facilitated by the synthesis process, which relies on the input signal ${\mathbf{X}}$. Following the spectral index and filter bank cases in Sec. 3.3, we define the Index NLSFs and Value NLSFs by
$\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{ind})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{ind})}_{{\bm{\Delta}}}({\mathbf{X}})))=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (4)
$\Theta_{\text{val}}({\bm{\Delta}},{\mathbf{X}})$ $=\mathcal{S}^{(\text{val})}_{{\bm{\Delta}},{\mathbf{X}}}(\Psi(\mathcal{A}^{(\text{val})}_{{\bm{\Delta}}}({\mathbf{X}})))=\sum_{j=1}^{K}\left[\Psi\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{K}\right]_{j}\ \frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e},$ (5)
where $\Psi:{\mathbb{R}}^{Jd}\rightarrow{\mathbb{R}}^{Jd}$ in $\Theta_{\text{ind}}$ and $\Psi:{\mathbb{R}}^{Kd}\rightarrow{\mathbb{R}}^{Kd}$ in $\Theta_{\text{val}}$. To adjust the feature output dimension at each node, we apply an MLP with shared weights to all nodes after the NLSF.
Graph-level NLSFs. We first introduce the Graph-level NLSFs that are fully spectral, where the NLSFs map a sequence of frequency coefficients to an output vector. Specifically, the Index-based and Value-based Graph-level NLSFs are given by
$\Phi_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)\text{\ \ and \ \ }\Phi_{\text{val}}({\bm{\Delta}},{\mathbf{X}})=\Psi\left(\left\lVert g_{i}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}\right).$ (6)
where $\Psi$ maps from ${\mathbb{R}}^{Jd}$ (resp. $\Phi_{\text{ind}}$) or ${\mathbb{R}}^{Kd}$ (resp. $\Phi_{\text{val}}$) to ${\mathbb{R}}^{d^{\prime}}$, and $d^{\prime}$ is the output dimension.
Pooling-NLSFs. We introduce another type of graph-level NLSFs by first representing each graph in a Node-level NLSFs as in Eq. (4) and Eq. (5). The final graph representation is obtained by applying a nonlinear activation function followed by a readout function to these node-level representations. We consider four commonly used pooling methods, including mean, sum, max, and $L_{p}$-norm pooling, as the readout function for each graph. We apply an MLP after readout function to obtain a $d^{\prime}$-dimensional graph-level representation. We term these graph-level NLSFs as Pooling-NLSFs.
3.5 Laplacian Attention NLSFs
To understand which Laplacian and parameterization (index v/s value) of the NLSF are preferable in different settings, we follow the random geometric graph analysis outlined in [Paolino et al., 2023]. Specifically, we consider a setting where random geometric graphs are sampled from a metric-probability spaces $\mathcal{S}$. In such a case, the graph Laplacian approximates continuous Laplacians on the metric spaces under some conditions. We aim for our NLSFs to produce approximately the same outcome for any two graphs sampled from the same underlying metric space $\mathcal{S}$, ensuring that the NLSF is transferable. In App. A.5, we show that if the nodes of the graph are sampled uniformly from $\mathcal{S}$, then using the graph Laplacian $\mathbf{L}$ in Index NLSFs yields a transferable method. Conversely, if the nodes of the graph are sampled non-uniformly, and any two balls of the same radius in $\mathcal{S}$ have the same volume, then utilizing the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs is a transferable method.
Given that real-life graphs often fall between these two boundary cases, we present an architecture that chooses between the Index NLSFs and Value NLSFs, as illustrated in Fig. 2. Specifically, a soft attention mechanism is employed to dynamically choose between the two parameterizations, given by
$\mathtt{att}\left(\Theta_{\text{ind}}(\mathbf{L},{\mathbf{X}}),\Theta_{\text{val}}(\mathbf{N},{\mathbf{X}})\right)=\alpha\Theta_{\text{ind}}(\mathbf{L},\mathbf{x})\|(1-\alpha)\Theta_{\text{val}}(\mathbf{N},{\mathbf{X}}),$
where $0\leq\alpha\leq 1$ is obtained using a softmax function to normalize the scores into attention weights, balancing each NLSFs’ contribution. It allows the model to adaptively select the most appropriate parameterization from the graph characteristics, promoting the transferability of the NLSFs.
Figure 2: Illustration of Laplacian attention NLSFs. An attention mechanism is applied to both Index NLSFs and Value NLSFs, enabling the adaptive selection of the most appropriate parameterization.
4 Theoretical Properties of Nonlinear Spectral Filters
We present the desired theoretical properties of our NLSFs at the node-level and graph-level.
4.1 Complexity of NLSFs
NLSFs are implemented by computing the eigenvectors of the GSO. Most existing spectral GNNs avoid direct eigendecomposition due to its perceived inefficiency. Instead, they use filters implemented by applying polynomials [Kipf and Welling, 2016, Defferrard et al., 2016a] or rational functions [Levie et al., 2018, Bianchi et al., 2021] to the GSO in the spatial domain. However, power iteration-based eigendecomposition algorithms, e.g., variants of the Lanczos method, can be highly efficient [Saad, 2011, Lanczos, 1950]. For matrices with $E$ non-zero entries, the computational complexity of one iteration for finding $J$ eigenvectors corresponding to the smallest or largest eigenvalues (called leading eigenvectors) is $O(JE)$. In practice, these eigendecomposition algorithms converge quickly due to their super-exponential convergence rate, often requiring only a few iterations.
This makes NLSFs applicable to node-level tasks on large sparse graphs, as they rely solely on the leading eigenvectors. In Sec. 4.4, we show that using the leading eigenvectors can approximate GSO well in the context of learning on graphs. For node-level tasks, such as semi-supervised node classification, the leading eigenvectors only need to be pre-computed once, with a complexity of $O(JE)$. During the learning phase, each step of the architecture search and hyperparameter optimization takes $O(NJd)$ complexity for analysis and synthesis, and $O(J^{2}d^{2})$ for the MLP in the spectral domain, which is faster than the complexity $O(Ed^{2})$ of message passing or standard spectral methods if $NJ<Ed$. Detailed empirical studies on runtime analysis are provided in the App. C.
For dense matrices, the computational complexity of a full eigendecomposition is $O(N^{b})$ per iteration, where $N^{b}$ is the complexity of matrix multiplication. This is practical for graph-level tasks on relatively small and dense graphs, which is typical for many graph classification datasets. In these cases, the eigendecomposition of all graphs in the dataset can be performed as a pre-computation step, significantly reducing the complexity during the learning phase.
4.2 Equivariance of Node-level NLSFs
We demonstrate the node-level equivariance of our NLSFs, ensuring that our method respects the functional shift symmetries. The proof is given in App. A.1.
Proposition 4.1.
Index NLSFs in Eq. (4) are equivariant to the graph functional shift $\mathcal{U}_{{\bm{\Delta}}}$, and Value NLSFs in Eq. (5) are equivariant to the relaxed graph functional shifts $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
4.3 Universal Approximation and Expressivity
In this subsection, we discuss the approximation power of NLSFs.
Node-Level Universal Approximation. We begin with a setting where a graph is given as a fixed domain, and the data distribution consists of multiple signals defined on this graph. An example of this setup is a spatiotemporal graph [Cini and Marisca, 2022], e.g., traffic networks, where a fixed sensor system defines a graph and the different signals represent the sensor readings collected at different times.
In App. A.2.1, we prove the following lemma, which shows that linear NLSFs exhaust the space of linear operators that commute with graph functional shifts.
Lemma 4.1.
Suppose that ${\scriptstyle\sum}_{j}\mathbf{P}_{j}={\scriptstyle\sum}_{j}g_{j}({\bm{\Delta}})=\mathbf{I}$. A linear operator ${\mathbb{R}}^{N\times d}\rightarrow{\mathbb{R}}^{N\times d}$ commute with $\mathcal{U}^{J}_{{\bm{\Delta}}}$ (resp. $\mathcal{U}_{{\bm{\Delta}}}^{g}$) iff it is a NLSF based on a linear function $\Psi$ in Eq. (4) (resp. Eq. (5)).
Lemma 4.1 shows a close relationship between functions that commute with functional shifts and those defined in the spectral domain. This motivates the following construction of a pseudo-metric on ${\mathbb{R}}^{N\times d}$. In the case of relaxed functional shifts, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We pull back the Euclidean metric to the spatial domain to define a signal pseudo-metric. Namely, for two signals ${\mathbf{X}}$ and ${\mathbf{X}}^{\prime}$, their distance is defined by
${\rm dist}_{{\bm{\Delta}}}\big{(}{\mathbf{X}},{\mathbf{X}}^{\prime}\big{)}:={\rm dist}_{\rm E}\big{(}\mathcal{A}({\bm{\Delta}},{\mathbf{X}}),\mathcal{A}({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})\big{)}.$
This pseudo metric can be made into a metric by considering each equivalence class of signals with zero distance as a single point in the space. As MLPs $\Psi$ can approximate any continuous function ${\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{K\times d}$ (the universal approximation theorem [Cybenko, 1989, Funahashi, 1989, Leshno et al., 1993]), node-level NLSFs can approximate any continuous function that maps (equivalence classes of) signals to (equivalence classes of) signals. For details, see App. A.2.2. A similar analysis applies to hard functional shifts.
Graph-Level Universal Approximation. The above analysis also motivates the construction of a graph-signal metric for graph-level tasks. For graphs with $d$-channel signals, we consider again the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We define the distance between any two graphs with GSOs and signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$ to be
${\rm dist}\big{(}({\bm{\Delta}},{\mathbf{X}}),({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})\big{)}:={\rm dist}_{\rm E}\big{(}\mathcal{A}({\bm{\Delta}},{\mathbf{X}}),\mathcal{A}({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})\big{)}.$
This definition can be extended into a metric by considering the space $\mathcal{G}$ of equivalence classes of graph-signals with distance 0. As before, this distance inherits the universal approximation properties of standard MLPs. Namely, any continuous function $\mathcal{G}\rightarrow{\mathbb{R}}^{d^{\prime}}$ with respect to ${\rm dist}$ can be approximated by NLSFs based on MLPs. Additional details are in App. A.2.3.
Graph-Level Expressivity of Pooling-NLSFs. In App. A.2.4, we show that Pooling-NLSFs are more expressive than graph-level NLSF when any $L_{p}$ norm is used in Eq. (4) and Eq. (5) with $p\neq 2$, both in the definition of the NLSF and as the pooling method. Specifically, for every graph-level NLSF, there is a Pooling-NLSF that coincides with it. Additionally, there are graph signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$ for which a Pooling-NLSF can attain different values, whereas any graph-level NLSF must attain the same value. Hence, Pooling-NLSFs have improved discriminative power compared to graph-level NLSFs. Indeed, as shown in Tab. LABEL:tab:graph_classification_accuracy_vertical, Pooling-NLSFs outperform Graph-level NLSFs in practice, which can be attributed to their increased expressivity.
4.4 Uniform Approximation of GSOs by Their Leading Eigenvectors
Since NLSFs on large graphs are based on the leading eigenvectors of ${\bm{\Delta}}$, we justify its low-rank approximation in the following. While approximating matrices with low-rank matrices might lead to a high error in the spectral and Frobenius norms, we show that such an approximation entails a uniformly small error in the cut norm. We define and interpret the cut norm in App. A.4.1, and explain why it is a natural graph similarity measure for graph machine learning.
The following theorem is a corollary of the Constructive Weak Regularity Lemma presented in [Finkelshtein et al., 2024]. Its proof is presented in App. A.4.
Theorem 4.1.
Let $\mathbf{M}$ be a symmetric matrix with entries bounded by $\left\lvert m_{i,j}\right\rvert\leq\alpha$, and let $J\in\mathbb{N}$. Suppose $m$ is sampled uniformly from $[J]$, and let $R\geq 1$ s.t. $J/R\in\mathbb{N}$. Let $\bm{\phi}_{1},\ldots,\bm{\phi}_{m}$ be the leading eigenvectors of $\mathbf{M}$, with eigenvalues $\mu_{1},\ldots,\mu_{m}$ ordered by their magnitudes $\left\lvert\mu_{1}\right\rvert\geq\ldots\geq\left\lvert\mu_{m}\right\rvert$. Define $\mathbf{C}=\sum_{k=1}^{m}\mu_{k}\bm{\phi}_{k}\bm{\phi}_{k}^{\top}$. Then, with probability $1-\frac{1}{R}$ (w.r.t. the choice of $m$),
$\|\mathbf{M}-\mathbf{C}\|_{\square}<\frac{3\alpha}{2}\sqrt{\frac{R}{J}}.$
Note that the bound in Thm. 4.1 is uniformly small, independently of $\mathbf{M}$ and its dimension $N$. This theorem justifies using the leading eigenvectors when working with the adjacency matrix as the GSO. For a justification when working with other GSOs see App. A.4.
5 Experimental Results
We evaluate the NLSFs on node and graph classification tasks. Additional implementation details are in App. B. Additional experiments, including runtime analysis and ablation studies, are in App. C.
5.1 Semi-Supervised Node Classification
Table 1: Semi-supervised node classification accuracy.
Cora Citeseer Pubmed Chameleon Squirrel Actor
GCN 81.92$\pm$0.9 70.73$\pm$1.1 80.14$\pm$0.6 43.64$\pm$1.9 33.26$\pm$0.8 27.63$\pm$1.7
GAT 83.64$\pm$0.7 71.32$\pm$1.3 79.45$\pm$0.7 42.19$\pm$1.3 28.21$\pm$0.9 29.46$\pm$0.9
SAGE 74.01$\pm$2.1 66.40$\pm$1.2 79.91$\pm$0.9 41.92$\pm$0.7 27.64$\pm$2.1 30.85$\pm$1.8
ChebNet 79.72$\pm$1.1 70.48$\pm$1.0 76.47$\pm$1.5 44.95$\pm$1.2 33.82$\pm$0.8 27.42$\pm$2.3
ChebNetII 83.95$\pm$0.8 71.76$\pm$1.2 81.38$\pm$1.3 46.37$\pm$3.1 34.40$\pm$1.1 33.48$\pm$1.2
CayleyNet 81.76$\pm$1.9 68.32$\pm$2.3 77.48$\pm$2.1 38.29$\pm$3.2 26.53$\pm$3.3 30.62$\pm$2.8
APPNP 83.19$\pm$0.8 71.93$\pm$0.8 82.69$\pm$1.4 37.43$\pm$1.9 25.68$\pm$1.3 35.98$\pm$1.3
GPRGNN 82.82$\pm$1.3 70.28$\pm$1.4 81.31$\pm$2.6 39.27$\pm$2.3 26.09$\pm$1.3 31.47$\pm$1.6
ARMA 81.64$\pm$1.2 69.91$\pm$1.6 79.24$\pm$0.5 39.40$\pm$1.8 27.42$\pm$0.7 30.42$\pm$2.6
$\mathtt{att}$-Node-level NLSFs 84.75$\pm$0.7 73.62$\pm$1.1 81.93$\pm$1.0 49.68$\pm$1.6 38.25$\pm$0.7 34.72$\pm$0.9
We first demonstrate the main advantage of the proposed Node-level NLSFs over existing GNNs with convolution design on semi-supervised node classification tasks. We test three citation networks [Sen et al., 2008, Yang et al., 2016]: Cora, Citeseer, and Pubmed. In addition, we explore three heterophilic graphs: Chameleon, Squirrel, and Actor [Rozemberczki et al., 2021, Tang et al., 2009]. For more comprehensive descriptions of these datasets, see App. B.
We compare the Node-level NLSFs using Laplacian attention with existing spectral GNNs for node-level predictions, including GCN [Kipf and Welling, 2016], ChebNet [Defferrard et al., 2016a], ChebNetII [He et al., 2022], CayleyNet [Levie et al., 2018], APPNP [Klicpera et al., 2019], GPRGNN [Chien et al., 2020], and ARMA [Bianchi et al., 2021]. We also consider GAT [Veličković et al., 2017] and SAGE [Hamilton et al., 2017]. For datasets splitting on citation graphs (Cora, Citeseer, and Pubmed), we apply the standard splits following [Yang et al., 2016], using 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For heterophilic graphs (Chameleon, Squirrel, and Actor), we use the sparse splitting as in [Chien et al., 2020, He et al., 2022], allocating 2.5% of samples for training, 2.5% for validation, and 95% for testing. We measure the classification quality by computing the average classification accuracy with a 95% confidence interval over 10 random splits. We utilize the source code released by the authors for the baseline algorithms and optimize their hyperparameters using Optuna [Akiba et al., 2019]. Each model’s hyperparameters are fine-tuned to achieve the highest possible accuracy. Detailed hyperparameter settings are provided in App. B.
Tab. LABEL:tab:node_classification_accuracy presents the node classification accuracy of our Node-level NLSFs using Laplacian attention and the various competing baselines. We see that $\mathtt{att}$-Node-level NLSFs outperform the competing models on the Cora, Citeseer, and Chameleon datasets. Notably, it shows remarkable performance on the densely connected Squirrel graph, outperforming the baselines by a large margin. This can be explained by the sparse version in Eq. (9) of Thm. 4.1, which shows that the denser the graphs is, the better its rank-$J$ approximation. For the Pubmed and Actor datasets, $\mathtt{att}$-Node-level NLSFs yield the second-best results, which are comparable to the best results obtained by APPNP.
5.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Table 2: Node classification performance on original and filtered Chameleon and Squirrel datasets.
Chameleon Squirrel
Original Filtered Original Filtered
ResNet+SGC 49.93$\pm$2.3 41.01$\pm$4.5 34.36$\pm$1.2 38.36$\pm$2.0
ResNet+adj 71.07$\pm$2.2 38.67$\pm$3.9 65.46$\pm$1.6 38.37$\pm$2.0
GCN 50.18$\pm$3.3 40.89$\pm$4.1 39.06$\pm$1.5 39.47$\pm$1.5
GPRGNN 47.26$\pm$1.7 39.93$\pm$3.3 33.39$\pm$2.1 38.95$\pm$2.0
FSGNN 77.85$\pm$0.5 40.61$\pm$3.0 68.93$\pm$1.7 35.92$\pm$1.3
GloGNN 70.04$\pm$2.1 25.90$\pm$3.6 61.21$\pm$2.0 35.11$\pm$1.2
FAGCN 64.23$\pm$2.0 41.90$\pm$2.7 47.63$\pm$1.9 41.08$\pm$2.3
$\mathtt{att}$-Node-level NLSFs 79.42$\pm$1.6 42.06$\pm$1.3 67.81$\pm$1.4 42.18$\pm$1.2
Recently, the work in [Platonov et al., 2023] identified the presence of many duplicate nodes across the train, validation, and test splits in the dense split setting of the Chameleon and Squirrel [Pei et al., 2020]. This results in train-test data leakage, causing GNNs to inadvertently fit the test splits during training, thereby making performance results on Chameleon and Squirrel less reliable. To further validate the performance of our Node-level NLSFs on these datasets in the dense split setting, we use both the original and filtered versions of Chameleon and Squirrel, which do not contain duplicate nodes, as suggested in [Platonov et al., 2023]. We use the same random splits as in [Platonov et al., 2023], dividing the datasets into 48% for training, 32% for validation, and 20% for testing.
Tab. 2 presents the classification performance comparison between the original and filtered Chameleon and Squirrel. The baseline results are taken from [Platonov et al., 2023], and we include the following competitive models: ResNet+SGC [Platonov et al., 2023], ResNet+adj [Platonov et al., 2023], GCN [Kipf and Welling, 2016], GPRGNN [Chien et al., 2020], FSGNN [Maurya et al., 2022], GloGNN [Li et al., 2022], and FAGCN [Bo et al., 2021]. The detailed comparisons are in App. C. We see in the table that the $\mathtt{att}$-Node-level NLSFs consistently outperform the competing baselines on both the original and filtered datasets. $\mathtt{att}$-Node-level NLSFs demonstrate less sensitivity to node duplicates and exhibit stronger generalization ability, further validating the reliability of the Chameleon and Squirrel datasets in the dense split setting. We note that compared to the dense split setting, the sparse split setting in Tab. LABEL:tab:node_classification_accuracy is more challenging, resulting in lower classification performance. A similar trend of significant performance difference between the two settings of Chameleon and Squirrel is also observed in [He et al., 2022].
5.3 Graph Classification
We further illustrate the power of NLSFs on eight graph classification benchmarks [Kersting et al., 2016]. Specifically, we consider five bioinformatics datasets [Borgwardt et al., 2005, Kriege and Mutzel, 2012, Shervashidze et al., 2011]: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, where MUTAG, PTC, and NCI1 are characterized by discrete node features, while ENZYMES and PROTEINS have continuous node features. Additionally, we examine three social network datasets: IMDB-B, IMDB-M, and COLLAB. The unattributed graphs are augmented by adding node degree features following [Errica et al., 2019]. For more details of these datasets, see App. B.
In graph classification tasks, a readout function is used to summarize node representations for each graph. The final graph-level representation is obtained by aggregating these node-level summaries and is then fed into an MLP with a (log)softmax layer to perform the graph classification task. We compare our NLSFs with two kernel-based approaches: GK [Shervashidze et al., 2009] and WL [Shervashidze et al., 2011], as well as nine GNNs: GCN [Kipf and Welling, 2016], GAT [Veličković et al., 2017], SAGE [Hamilton et al., 2017], ChebNet [Defferrard et al., 2016a], ChebNetII [He et al., 2022], CayleyNet [Levie et al., 2018], APPNP [Klicpera et al., 2019], GPRGNN [Chien et al., 2020], and ARMA [Bianchi et al., 2021]. Additionally, we consider the hierarchical graph pooling model DiffPool [Ying et al., 2018]. For dataset splitting, we apply the random split following [Velickovic et al., 2019, Ying et al., 2018, Ma et al., 2019, Zhu et al., 2021b], using 80% for training, 10% for validation, and 10% for testing. This random splitting process is repeated 10 times, and we report the average performance along with the standard deviation. For the baseline algorithms, we use the source code released by the authors and optimize their hyperparameters using Optuna [Akiba et al., 2019]. We fine-tune the hyperparameters of each model to achieve the highest possible accuracy. Detailed hyperparameter settings for both the baselines and our method are provided in App. B.
Table 3: Graph classification accuracy.
MUTAG PTC ENZYMES PROTEINS NCI1 IMDB-B IMDB-M COLLAB
GK 76.38$\pm$2.9 52.13$\pm$1.2 30.07$\pm$6.1 72.33$\pm$4.5 62.62$\pm$2.2 67.63$\pm$1.0 44.19$\pm$0.9 72.32$\pm$3.7
WL 78.04$\pm$1.8 52.44$\pm$1.3 51.29$\pm$5.9 76.20$\pm$4.1 76.45$\pm$2.3 71.42$\pm$3.2 46.63$\pm$1.3 76.23$\pm$2.2
GCN 81.63$\pm$3.1 60.22$\pm$1.9 43.66$\pm$3.4 75.17$\pm$3.7 76.29$\pm$1.8 72.96$\pm$1.3 50.28$\pm$3.2 79.98$\pm$1.9
GAT 83.17$\pm$4.4 62.31$\pm$1.4 39.83$\pm$3.7 74.72$\pm$4.1 74.01$\pm$4.3 70.62$\pm$2.5 45.67$\pm$2.7 74.27$\pm$2.5
SAGE 75.18$\pm$4.7 61.33$\pm$1.1 37.99$\pm$3.7 74.01$\pm$4.3 74.90$\pm$1.7 68.38$\pm$2.6 46.94$\pm$2.9 73.61$\pm$2.1
DiffPool 82.40$\pm$1.4 56.43$\pm$2.9 60.13$\pm$3.2 79.47$\pm$3.1 77.18$\pm$0.7 71.09$\pm$1.6 50.43$\pm$1.5 80.16$\pm$1.8
ChebNet 82.15$\pm$1.6 64.06$\pm$1.2 50.42$\pm$1.4 74.28$\pm$0.9 76.98$\pm$0.7 73.14$\pm$1.1 49.82$\pm$1.6 77.40$\pm$1.6
ChebNetII 84.17$\pm$3.1 70.03$\pm$2.8 64.29$\pm$2.9 78.31$\pm$4.1 81.14$\pm$3.6 77.09$\pm$3.9 52.69$\pm$2.7 80.06$\pm$3.3
CayleyNet 83.06$\pm$4.2 62.73$\pm$5.1 42.28$\pm$5.3 74.12$\pm$4.7 77.21$\pm$4.5 71.45$\pm$4.8 51.89$\pm$3.9 76.33$\pm$5.8
APPNP 84.45$\pm$4.4 65.26$\pm$6.9 49.68$\pm$3.9 77.26$\pm$4.8 73.24$\pm$7.3 72.94$\pm$2.3 44.36$\pm$1.9 73.85$\pm$3.3
GPRGNN 80.26$\pm$2.0 58.41$\pm$1.4 45.29$\pm$1.7 73.90$\pm$2.5 73.12$\pm$4.0 69.17$\pm$2.6 47.07$\pm$2.8 77.93$\pm$1.9
ARMA 83.21$\pm$2.7 69.23$\pm$2.2 61.21$\pm$3.4 76.62$\pm$3.6 79.51$\pm$2.9 73.27$\pm$2.7 53.60$\pm$1.9 78.34$\pm$1.1
$\mathtt{att}$-Graph-level NLSFs 84.13$\pm$1.5 68.17$\pm$1.0 65.94$\pm$1.6 82.69$\pm$1.9 80.51$\pm$1.2 74.26$\pm$1.8 52.49$\pm$0.7 79.06$\pm$1.2
$\mathtt{att}$-Pooling-NLSFs 86.89$\pm$1.2 71.02$\pm$1.3 69.94$\pm$1.0 84.89$\pm$0.9 80.95$\pm$1.4 76.78$\pm$1.9 55.28$\pm$1.7 82.19$\pm$1.3
Tab. LABEL:tab:graph_classification_accuracy_vertical presents the graph classification accuracy. Notably, the $\mathtt{att}$-Graph-level NLSFs (i.e., NLSFs without the synthesis and pooling process) perform the second best on the ENZYMES and PROTEINS. Additionally, $\mathtt{att}$-Pooling-NLSFs consistently outperform $\mathtt{att}$-Graph-level NLSFs, indicating that the node features learned in our Node-level NLSFs representation are more expressive, corroborating our theoretical findings in Sec. 4.3. Furthermore, our $\mathtt{att}$-Pooling-NLSFs consistently outperform all baselines on the MUTAG, PTC, ENZYMES, PROTEINS, IMDB-M, and COLLAB datasets. For NCI1 and IMDB-B, $\mathtt{att}$-Pooling-NLSFs rank second and are comparable to ChebNetII.
6 Summary
We presented an approach for defining non-linear filters in the spectral domain of graphs in a transferable and equivariant way. Transferability between different graphs is achieved by using the input signal as a basis for the synthesis operator, making the NLSF depend only on the eigenspaces of the GSO and not on an arbitrary choice of the eigenvectors. While different graph-signals may be of different sizes, the spectral domain is a fixed Euclidean space independent of the size and topology of the graph. Hence, our spectral approach represents graphs as vectors. We note that standard spectral methods do not have this property since the coefficients of the signal in the frequency domain depend on an arbitrary choice of eigenvectors, while our representation depends only on the eigenspaces. We analyzed the universal approximation and expressivity power of NLSFs through metrics that are pulled back from this Euclidean vector space. From a geometric point of view, our NLSFs are motivated by respecting graph functional shift symmetries, making them related to Euclidean CNNs.
Limitation and Future Work. One limitation of NLSFs is that, when deployed on large graphs, they only depend on the leading eigenvectors of the GSO. In this paper, we considered the low-frequency eigenvectors of graph Laplacians. However, important information can also lie within the high frequencies or in any other band. In future work, we plan to explore NLSFs that are sensitive to eigenvalues that lie within a set of bands of interest, which can be adaptive to the graph.
Acknowledgments
The work of YEL and RT was supported by the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 802735-ERC-DIFFOP. The work of RL was supported by the Israel Science Foundation grant #1937/23: Analysis of graph deep learning using graphon theory.
Appendix A Theoretical Analysis
We note that the numbering of the statements corresponds to the numbering used in the paper, and we have also included several separately numbered propositions and lemmas that are used in supporting the proofs presented. We restate the claim of each statement for convenience.
A.1 Equivariance of NLSFs
We start with two lemmas that characterize the graph functional shifts. The lemmas directly follow the fact that two normal operators commute iff the projections upon their eigenspaces commute.
Projection to Eigenspaces Case.
Denote by $\mathcal{R}$ the span of $\bigcup_{j=1}^{J}\mathbf{P}_{j}{\mathbb{R}}^{N}$. Denote by $\mathcal{R}^{\perp}$ the orthonormal complement of $\mathcal{R}$. Note that
${\mathbb{R}}^{N}=\left(\prod_{j=1}^{J}\mathbf{P}_{j}{\mathbb{R}}^{N}\right)\times\mathcal{R}^{\perp},$
where $\prod$ and $\times$ denote direct products of linear spaces. Denote by $\oplus$ the direct sum of operators.
Lemma A.1.
$\mathbf{U}$ is in $\mathcal{U}_{{\bm{\Delta}}}$ iff it has the form $\mathbf{U}=(\bigoplus_{j=1}^{J}\mathbf{U}_{j})\oplus\mathbf{U}^{\perp}$, where $\mathbf{U}_{j}$ is a unitary operator in $\mathbf{P}_{j}{\mathbb{R}}^{N}$, and $\mathbf{U}^{\perp}$ a unitary operator in $\mathcal{R}^{\perp}$.
Projection to Bands Case.
Denote by $\mathcal{R}_{g}$ the span of $\bigcup_{j=1}^{K}g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$. Denote by $\mathcal{R}^{\perp}_{g}$ the orthonormal complement of $\mathcal{R}_{g}$ Note that
${\mathbb{R}}^{N}=\left(\prod_{j=1}^{K}g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}\right)\times\mathcal{R}^{\perp}_{g}.$
Lemma A.2.
$\mathbf{U}$ is in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ iff it has the form $\mathbf{U}=(\bigoplus_{j=1}^{K}\mathbf{U}_{j})\oplus\mathbf{U}^{\perp}_{g}$, where $\mathbf{U}_{j}$ is a unitary operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$, and $\mathbf{U}^{\perp}$ a unitary operator in $\mathcal{R}^{\perp}_{g}$.
A.1.1 Proof of Propositions 4.1
Proposition 4.1.
Index NLSFs in Eq. (4) are equivariant to the graph functional shift $\mathcal{U}_{{\bm{\Delta}}}$, and Value NLSFs in Eq. (5) are equivariant to the relaxed graph functional shifts $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
Proof.
We start with Index-NLSFs. Consider the Index NLSFs defined as in Eq. (4)
$\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}})=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
We need to show that for any unitary operator $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$,
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\mathbf{U}\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}}).$
Consider $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$ and apply it to the graph signal $\mathbf{X}$, the Index NLSFs with the transformed input is given by
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
Since $\mathbf{U}\in\mathcal{U}_{{\bm{\Delta}}}$, it commutes with $\mathbf{P}_{j}$ for all $j$, i.e., $\mathbf{U}\mathbf{P}_{j}=\mathbf{P}_{j}\mathbf{U}$. Using this commutation property, we can rewrite the norm and the projections:
$\left\lVert\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}\right\rVert_{\rm sig}=\left\lVert\mathbf{U}\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}.$
In addition, since $\mathbf{U}$ is a unitary matrix, it preserves the norm. Therefore, we have
$\left\lVert\mathbf{U}\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}=\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}.$
Substitute these expressions back into the definition of the Index NLSFs:
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}\mathbf{U}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
Notice that $\mathbf{U}$ appears linearly in the numerator of the fraction. We can factor it out:
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\mathbf{U}\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
The expression inside the summation is exactly the original Index NLSFs applied to
$\Theta_{\text{ind}}({\bm{\Delta}},\mathbf{U}{\mathbf{X}})=\mathbf{U}\Theta_{\text{ind}}({\bm{\Delta}},{\mathbf{X}}).$
Therefore, we have shown that applying the unitary operator $\mathbf{U}$ to the graph signal ${\mathbf{X}}$ results in the Index NLSFs being transformed by the same unitary operator $\mathbf{U}$, proving the equivariance property.
The proof for value-NLSF follows the same steps. ∎
A.2 Expressivity and Universal Approximation
In this section, we focus on value parameterized NLSFs. The analysis for index-NLSF is equivalent.
A.2.1 Proof of Lemma 4.1 - Linear Node-level NLSF Exhaust the Linear Operators that Commute with Functional Shifts
Lemma 4.1
Suppose that $\sum_{j}\mathbf{P}_{j}=\sum_{j}g_{j}({\bm{\Delta}})=\mathbf{I}$. A linear operator ${\mathbb{R}}^{N\times d}\rightarrow{\mathbb{R}}^{N\times d}$ commute with $\mathcal{U}^{J}_{{\bm{\Delta}}}$ (resp. $\mathcal{U}_{{\bm{\Delta}}}^{g}$) iff it is a NLSF based on a linear function $\Psi$ in Eq. (4) (resp. Eq. (5)).
Proof.
Consider a setting where a single graph is given as a fixed domain, and the data distribution consists of many signals defined on this graph. For simplicity, we restrict the analysis to the case of 1D signals ($d=1$). The extension to a general dimension $d$ is natural.
We next show that node-level linear NLSFs exhaust the space of linear operators that commute with graph functional shifts (on the fixed graph).
By Lemma A.2, any unitary operator $\mathbf{U}$ in $\mathcal{U}^{g}_{{\bm{\Delta}}}$ has the form
$\mathbf{U}=\left(\bigoplus_{j=1}^{K}\mathbf{U}_{j}\right)\oplus\mathbf{U}^{\perp},$
where $\mathbf{U}_{j}$ is a unitary operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$, and $\mathbf{U}^{\perp}$ a unitary operator in $\mathcal{R}^{\perp}_{g}$.
Hence, since the unitary representation $\mathbf{U}\mapsto\mathbf{U}$ of the group of unitary operators in ${\mathbb{R}}^{N}$ is irreducible, by Schur’s lemma [Serre et al., 1977] any linear operator $\mathbf{B}$ that commutes with all operators of $\mathcal{U}^{g}_{{\bm{\Delta}}}$ must be a scalar times the identity when projected to $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$ and to $\mathcal{R}_{g}^{\perp}$. Namely, $\mathbf{B}$ has the form
$\mathbf{B}=\left(\bigoplus_{j=1}^{K}(b_{j}\mathbf{I}_{j})\right)\oplus(b^{\perp})\mathbf{I}^{\perp},$
where $\mathbf{I}_{j}$ is the identity operator in $g_{j}({\bm{\Delta}}){\mathbb{R}}^{N}$, $\mathbf{I}^{\perp}$ is the identity operator in $\mathcal{R}^{\perp}_{g}$, and $b_{j}$ and $b^{\perp}$ are scalars. This means that linear node-level NLSFs exhaust the space of linear operators that commute with $\mathcal{U}^{g}_{{\bm{\Delta}}}$.
The case of hard graph functional shifts is treated similarly. ∎
NLSF with High Frequency Content.
We note that the condition $\sum_{j}\mathbf{P}_{j}=\sum_{j}g_{j}({\bm{\Delta}})=\mathbf{I}$ can be omitted from Lemma 4.1 if we extend the definition of NLSF by adding one more term in the spectrum as follows. In the spectral index case, concatenate to the analysis operator the coefficient $\left\lVert\mathbf{X}-\sum_{j}\mathbf{P}_{j}\mathbf{X}\right\rVert$, and in the value case, concatenate $\left\lVert\mathbf{X}-\sum_{j}g_{j}({\bm{\Delta}})\mathbf{X}\right\rVert$. Extend accordingly the definition of synthesis and of NLSFs to also depend on this term. We do not include this spectral term in practice as it does not lead to improved performance.
A.2.2 Node-Level Universal Approximation
The above analysis motivates the following construction of a metric on ${\mathbb{R}}^{N}$. Given the filter bank $\{g_{j}\}_{j=1}^{K}$, on graph with $d$-dimensional signals, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We pull back the Euclidean metric to the spatial domain to define a graph-signal metric. Namely, for two signals ${\mathbf{X}}$ and ${\mathbf{X}}^{\prime}$, their distance is defined to be
${\rm dist}_{{\bm{\Delta}}}\big{(}{\mathbf{X}},{\mathbf{X}}^{\prime}\big{)}:={\rm dist}_{\rm E}\big{(}\mathcal{A}({\bm{\Delta}},{\mathbf{X}}),\mathcal{A}({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})\big{)}.$
[…]
Now, since MLPs $\Psi$ can approximate any continuous function ${\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{K\times d^{\prime}}$ (by the universal approximation theorem), and by the fact that $\mathcal{A}_{{\bm{\Delta}}}/{\rm dist}_{{\bm{\Delta}}}$ is an isometry, node-level NLSFs based on MLPs in the spectral domain can approximate any continuous function from signals with $d$ features to signal with $d^{\prime}$ features $(\mathcal{N}_{{\bm{\Delta}}})^{d}\rightarrow(\mathcal{N}_{{\bm{\Delta}}})^{d^{\prime}}$.
A.2.3 Graph-Level Universal approximation
The above analysis also motivates the construction of a graph-signal metric for graph-level tasks. For graphs with $d$-channel signals, we define the standard Euclidean metric ${\rm dist}_{\rm E}$ in the spectral domain ${\mathbb{R}}^{K\times d}$. We pull back the Euclidean metric to the spatial domain to define a graph-signal metric. Namely, for two graphs with Laplacians and signals $({\bm{\Delta}},{\mathbf{X}})$ and $({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})$, their distance is defined to be
${\rm dist}\big{(}({\bm{\Delta}},{\mathbf{X}}),({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})\big{)}:={\rm dist}_{\rm E}\big{(}\mathcal{A}({\bm{\Delta}},{\mathbf{X}}),\mathcal{A}({\bm{\Delta}}^{\prime},{\mathbf{X}}^{\prime})\big{)}$
This defines a pseudometric on the space of graph-signals. To obtain a metric, we consider equivalence classes of graph-signals with zero distance as a single point. Namely, we define the space $\mathcal{G}$ to be the space of graph-signal modulo ${\rm dist}$. In $\mathcal{G}$, the function ${\rm dist}$ becomes a metric, and $\mathcal{A}$ an isometry.
By the isometry property, $\mathcal{G}$ inherits any approximation property from ${\mathbb{R}}^{K\times d}$. For example, since MLPs can approximate any continuous function ${\mathbb{R}}^{K\times d}\rightarrow{\mathbb{R}}^{d^{\prime}}$, the space of NLSFs based on MLPs $\Psi$ has a universality property: any continuous function $\mathcal{G}\rightarrow{\mathbb{R}}^{d^{\prime}}$ with respect to ${\rm dist}$ can be approximated by a NLSF based on an MLP.
A.2.4 Graph-Level Expressivity of Pooling-NLSFs
We now show that pooling-NLSF are more expressive than graph-level NLSF if the norm in Eq. (4) and Eq. (5) is $L_{p}$ with $p\neq 2$.
First, we show that there are graph-signals that graph-level-NLSFs cannot separate and Pooling-NLSFs can. Consider an index NLSF with norm $L_{1}$ normalize by $1/N$. Namely, for $\mathbf{a}\in{\mathbb{R}}^{N}$,
$\left\lVert\mathbf{a}\right\rVert_{1}=\frac{1}{N}\sum_{n=1}^{N}\left\lvert a_{n}\right\rvert.$
The general case is similar.
For the two graphs, take the graph Laplacian
$\left(\begin{array}[]{rr}1&-1\\ -1&1\end{array}\right)$
with eigenvalues $0,1$, and corresponding eigenvectors $(1,1)$ and $(1,-1)$. Take the graph Laplacian
$\left(\begin{array}[]{rrr}1&-1&0\\ -1&2&-1\\ 0&-1&1\end{array}\right)$
with eigenvalues $0,1,3$. The first two eigenvectors are $(1,1,1)$ and $(1,0,-1)$ respectively.
Consider an Index-NLSF based on two eigenprojections $\mathbf{P}_{1},\mathbf{P}_{2}$. As the signal of the first graph take $(1,1)+(1,-1)$, and for the second graph take $(1,1,1)+\frac{3}{2}(1,0,-1)$. Both graph-signals have the same spectral coefficients $(1,1)$, so graph-level NLSF cannot separate them. Suppose that the NLSF is
$\Theta_{\text{val}}({\bm{\Delta}},{\mathbf{X}})=(\left\lVert\mathbf{P}_{1}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e)\frac{\mathbf{P}_{1}{\mathbf{X}}}{\left\lVert\mathbf{P}_{1}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\ +\ (\left\lVert\mathbf{P}_{2}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e)\frac{\mathbf{P}_{2}{\mathbf{X}}}{\left\lVert\mathbf{P}_{2}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
The outcome of the corresponding Pooling NLSF on the two graphs is
$1=\left\lVert(1+1,1-1)\right\rVert_{1}\neq\left\lVert(1+3/2,1,1-3/2)\right\rVert_{1}=\frac{4}{3}.$
Hence, Pooling-NLSFs separate these inputs, while graph-level-NLSFs do not.
Next, we show that Pooling-NLSFs are at least as expressive as graph-level NLSFs. In particular, any graph-level NLSF can be expressed as a pooling NLSF.
Let $\Psi$ be a graph-level NLSF. Define the node-level NLSF that chooses one spectral index $j$ with a nonzero value (e.g., the band with the largest coefficient) and projects upon it the value $\Psi({\bm{\Delta}},{\mathbf{X}})(\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e)/\left\lVert\mathbf{P}_{j}X\right\rVert_{\rm sig}$. Hence, before pooling, the NLSF gives
$\Theta({\bm{\Delta}},{\mathbf{X}})=\Psi({\bm{\Delta}},{\mathbf{X}})(\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e)\frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}/\left\lVert\mathbf{P}_{j}X\right\rVert_{\rm sig},$
where $j$ depends on the spectral coefficients (e.g., it is the index of the largest spectral coefficient). Hence, after pooling, the Pooling NLSF returns $\Psi({\bm{\Delta}},{\mathbf{X}})$, which coincides with the output of the graph-level NLSF.
Now, let us show that for $p=2$, they have the same expressivity. For any NLSF,
$\Theta({\bm{\Delta}},{\mathbf{X}})$ $=\left\lVert\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}\ \frac{\mathbf{P}_{j}{\mathbf{X}}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}\right\rVert^{2}_{\rm sig}$
$=\sum_{j=1}^{J}\left[\Psi\left(\left\lVert\mathbf{P}_{i}{\mathbf{X}}\right\rVert_{\rm sig}\right)_{i=1}^{J}\right]_{j}^{2}\ \frac{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert^{2}_{\rm sig}}{\left\lVert\mathbf{P}_{j}{\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}.$
This is a generic fully spectral NLSF.
A.3 Stable Invertibility of Synthesis
In the fixed graph setting, given any signal ${\mathbf{X}}$ such that $g_{j}({\bm{\Delta}}){\mathbf{X}}\neq 0$ for every $j\in[K]$, the synthesis operator $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}$ is stably invertible. Indeed, the right singular vectors of $\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}$ are the standard basis elements $\mathbf{e}_{j}$ in ${\mathbb{R}}^{K}$, the left singular vectors are
$\mathbf{Y}_{j}=\frac{g_{j}({\bm{\Delta}}){\mathbf{X}}}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}}\in{\mathbb{R}}^{N},$
and the singular values are
$\sigma_{j}=\frac{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{\rm sig}^{a}+e}{\left\lVert g_{j}({\bm{\Delta}}){\mathbf{X}}\right\rVert_{2}},$
for $j\in[K]$. Hence, we have
$\left\lVert\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}\right\rVert_{2}=\max_{j}\sigma_{j}$
and
$\left\lVert\mathcal{S}_{{\bm{\Delta}},{\mathbf{X}}}^{-1}\right\rVert_{2}=\frac{1}{\min_{j}\sigma_{j}}$
Suppose that $a=1$ and $e=0$. In this case, $\mathcal{S}$ is an isometry from the spectral domain to a subspace of the signal space ${\mathbb{R}}^{N}$. Analysis is the adjoint of synthesis.
This analysis can be extended to the index parametrization case.
A.4 Uniform Approximation of Graphs with $J$ Eigenvectors
In this section, we develop the setting under which the low-rank approximation of GSOs with their leading eigenvectors can be interpreted as a uniform approximation (Sec. 4.4).
A.4.1 Cut Norm
The cut norm of $\mathbf{M}\in{\mathbb{R}}^{N\times N}$ is defined to be
$\left\lVert\mathbf{M}\right\rVert_{\square}:=\frac{1}{N^{2}}\sup_{S,T\subset[N]}\Big{|}\sum_{i\in S}\sum_{j\in T}m_{i,j}\Big{|}.$ (7)
The distance between two matrices in cut norm is defined to be $\left\lVert\mathbf{M}-\mathbf{M}^{\prime}\right\rVert_{\square}$.
The cut norm has the following interpretation, which has precise formulation in terms of the weak regularity lemma [Frieze and Kannan, 1999, Lovász and Szegedy, 2007]. Any pair of (deterministic) graphs are close to each other in cut norm if and only if they can be described as pseudo-random graphs sampled from the same stochastic block model. Hence, the cut norm is a meaningful notion of graph similarity for practical graph machine learning, where graphs are noisy and can represent the same underlying phenomenon even if they have different sizes and topologies. In addition, the distance between non-isomorphic graphons is always positive in cut norm [Lovász, 2012]. In this context, the work in [Levie, 2023] showed that GNNs with normalized sum aggregation cannot separate graphs that have zero distance in the cut norm. This means that the cut norm is sufficiently discriminative for practical machine learning on graphs.
A.4.2 The Constructive Weak Regularity Lemma in Hilbert Spaces
The following lemma, called the constructive weak regularity lemma in Hilbert spaces, was proven in [Finkelshtein et al., 2024]. It is an extension of the classical respective result from [Lovász and Szegedy, 2007].
We define the Frobenius norm normalized by $1/N$ as
$\left\lVert\mathbf{M}\right\rVert_{\rm F}=\sqrt{\frac{1}{N}\sum_{i,j=1}^{N}\left\lvert m_{i,j}\right\rvert^{2}}.$
Lemma A.3.
[[Finkelshtein et al., 2024]] Let $\{\mathcal{K}_{j}\}_{j\in\mathbb{N}}$ be a sequence of nonempty subsets of a real Hilbert space $\mathcal{H}$ and let $\delta\geq 0$. Let $J>0$, let $R\geq 1$ such that $J/R\in\mathbb{N}$, and let $g\in\mathcal{H}$. Let $m$ be randomly uniformly sampled from $[J]$. Then, in probability $1-\frac{1}{R}$ (with respect to the choice of $m$), any vector of the form
$g^{*}=\sum_{j=1}^{m}\gamma_{j}f_{j}\quad\text{\ \ such that\ \ }\quad\bm{\gamma}=(\gamma_{j})_{j=1}^{m}\in\mathbb{R}^{m}\quad\text{and}\quad\mathbf{f}=(f_{j})_{j=1}^{m}\in\mathcal{K}_{1}\times\ldots\times\mathcal{K}_{m}$
that gives a close-to-best Hilbert space approximation of $g$ in the sense that
$\|g-g^{*}\|\leq(1+\delta)\inf_{\bm{\gamma},\mathbf{f}}\|g-\sum_{i=1}^{m}\gamma_{i}f_{i}\|,$ (8)
where the infimum is over $\bm{\gamma}\in{\mathbb{R}}^{m}$ and $\mathbf{f}\in\mathcal{K}_{1}\times\ldots\times\mathcal{K}_{m}$, also satisfies
$\forall w\in\mathcal{K}_{m+1},\quad\left\lvert\langle w,g-g^{*}\rangle\right\rvert\leq\left\lVert w\right\rVert\left\lVert g\right\rVert\sqrt{\frac{R}{J}+\delta}.$
A.4.3 Proof of Theorem 4.1
Theorem 4.1.
Let $\mathbf{M}$ be a symmetric matrix with entries bounded by $\left\lvert m_{i,j}\right\rvert\leq\alpha$, and let $J\in\mathbb{N}$. Suppose $m$ is sampled uniformly from $[J]$, and let $R\geq 1$ s.t. $J/R\in\mathbb{N}$. Consider $\bm{\phi}_{1},\ldots,\bm{\phi}_{m}$ as the leading eigenvectors of $\mathbf{M}$, with eigenvalues $\mu_{1},\ldots,\mu_{m}$ ordered by their magnitudes $\left\lvert\mu_{1}\right\rvert\geq\ldots\geq\left\lvert\mu_{m}\right\rvert$. Define $\mathbf{C}=\sum_{k=1}^{m}\mu_{k}\bm{\phi}_{k}\bm{\phi}_{k}^{\top}$. Then, with probability $1-\frac{1}{R}$ (w.r.t. the choice of $m$),
$\|\mathbf{M}-\mathbf{C}\|_{\square}<\frac{3\alpha}{2}\sqrt{\frac{R}{J}}.$
Proof.
Let us use Lemma A.3, with $\mathcal{H}={\mathbb{R}}^{N\times N}$, and $\mathcal{K}_{j}=\mathcal{K}$ the set of symmetric rank one matrices of the form $\mathbf{v}\mathbf{v}^{\top}$ where $\mathbf{v}\in{\mathbb{R}}^{N}$ is a column vector. Denote by $\mathcal{Y}_{m}$ the space of linear combinations of $m$ elements of $\mathcal{K}$, which is the space of symmetric matrices of rank bounded by $m$. For the Hilbert space norm, we take the Frobenius norm. In the setting of the lemma, we take $g=\mathbf{M}$, and $g^{*}\in\mathcal{Y}_{m}$, and $\delta=0$. By the lemma, with probability $1-1/R$, any Frobenius minimizer $\mathbf{C}$, namely, that satisfies $\left\lVert\mathbf{M}-\mathbf{C}\right\rVert_{\rm F}=\min_{\mathbf{C}^{\prime}\in\mathcal{Y}_{m}}\left\lVert\mathbf{M}-\mathbf{C}^{\prime}\right\rVert_{\rm F}$, also satisfies
$\left\langle\mathbf{Y},\mathbf{M}-\mathbf{C}\right\rangle\leq\left\lVert\mathbf{Y}\right\rVert_{\rm F}\left\lVert\mathbf{M}\right\rVert_{\rm F}\sqrt{\frac{R}{J}}\leq\alpha\left\lVert\mathbf{Y}\right\rVert_{\rm F}\sqrt{\frac{R}{J}}$
for every $\mathbf{Y}\in\mathcal{Y}_{m}$. Hence, for every choice of subset $S,T\subset[N]$, we have
$\left\lvert\sum_{i\in S}\sum_{j\in T}(m_{i,j}-c_{i,j})\right\rvert$
$=$ $\frac{1}{2}\left\lvert\sum_{i\in S}\sum_{j\in T}(m_{i,j}-c_{i,j})+\sum_{i\in T}\sum_{j\in S}(m_{i,j}-c_{i,j})\right\rvert$
$=$ $\frac{1}{2}\left\lvert\sum_{i\in S\cup T}\sum_{j\in S\cup T}(m_{i,j}-c_{i,j})\ -\ \sum_{i\in S}\sum_{j\in S}(m_{i,j}-c_{i,j})\ -\ \sum_{i\in T}\sum_{j\in T}(m_{i,j}-c_{i,j})\right\rvert$
$\leq$ $\left\lVert\frac{1}{2}\left\langle\bm{\text{1}}_{S\cup T}\bm{\text{1}}_{S\cup T}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}+\left\lVert\frac{1}{2}\left\langle\bm{\text{1}}_{S}\bm{\text{1}}_{S}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}+\left\lVert\frac{1}{2}\left\langle\bm{\text{1}}_{T}\bm{\text{1}}_{T}^{\top},\mathbf{M}-\mathbf{C}\right\rangle\right\rVert_{\rm F}$
$\leq$ $\frac{3\alpha}{2}\sqrt{\frac{R}{J}},$
where for a set $S\subset[N]$, denote by $\bm{\text{1}}_{S}\in{\mathbb{R}}^{N}$ the column vector with $1$ at coordinates in $S$ and 00 otherwise.
Hence, we also have
$\left\lVert\mathbf{M}-\mathbf{C}\right\rVert_{\square}\leq\frac{3\alpha}{2}\sqrt{\frac{R}{J}}$
Lastly, note that by the best rank-$m$ approximation theorem (Eckart–Young–Mirsky Theorem [Trefethen and Bau, 2022, Thm. 5.9]), any Frobenius minimizer $\mathbf{C}$ is the projection upon the $m$ leading eigenvectors of $\mathbf{M}$ (or some choice of these eigenvectors in case of multiplicity higher than 1).
Now, one can use the adjacency matrix $\mathbf{A}$ as $\mathbf{M}$ in Thm. 4.1. When working with sparse matrices of $E\ll N^{2}$ edges, to achieve a meaningful scaling of cut distance, we re-normalize the cut norm and define
$\left\lVert\mathbf{M}\right\rVert^{(E)}_{\square}=\frac{N^{2}}{E}\left\lVert\mathbf{M}\right\rVert_{\square}.$
With this norm, Thm. 4.1 gives
$\|\mathbf{M}-\mathbf{C}\|^{(E)}_{\square}<\frac{3\alpha N^{2}}{2E}\sqrt{\frac{R}{J}}.$ (9)
While this bound is not uniformly small in $N$, it is still independent of $\mathbf{M}$. In contrast, the error bounds for spectral and Frobenius norms do depend on the specific properties of $\mathbf{M}$.
Now, if we want to apply Thm. 4.1 to other GSOs ${\bm{\Delta}}$, we need to make some assumptions. Note that when the GSO is a Laplacian, we take as the leading eigenvectors the ones correspoding to the smallest eigenvalues, not the largest ones. To make the theorem applicable, we need to reorder the eigenvectors of ${\bm{\Delta}}$. This can be achieved by applying a decreasing function $h$ to ${\bm{\Delta}}$, such as $h({\bm{\Delta}})={\bm{\Delta}}^{-1}$. The role of $h$ is to amplify the eigenspaces of ${\bm{\Delta}}$ in which most of the energy of signals interest (the ones that often appear in the dataset) is concentrated. Under the assumption that $g({\bm{\Delta}})$ has entries bounded by some not-too-high $\alpha>0$, one can now justify approximating GSOs by low-rank approximations based on the smallest eigenvectors.
A.5 NLSFs on Random Geometric Graphs
In this section we consider the $L_{2}[N]$ norm normalized by $1/N^{1/2}$, namely,
$\left\lVert\mathbf{v}\right\rVert_{2}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}\left\lvert v_{n}\right\rvert^{2}}.$
We follow a similar analysis to [Paolino et al., 2023], mostly skipping the rigorous Monte-Carlo error rate proofs, as these are standard. Let $\mathcal{S}$ be a metric space with metric ${\rm d}$ and with a Borel probability measure, that we call the uniform measure $dx$. Let $r>0$, and denote by $B_{r}(x)$ the ball of radius $r$ about $x\in\mathcal{S}$. Let $V(x)$ be the volume of $B_{r}(x)$ with respect to the uniform measure. We consider an underlying Laplacian on the space $\mathcal{S}$ defined on signals $f:\mathcal{S}\rightarrow\mathbb{R}$ by
[…]
where we assume $C=1$ in the analysis below without loss of generality. Note that $\mathcal{L}$ is compact, so it has a discrete spectrum.
[…]
We consider the leading $M$ eigenvectors of $\mathbf{L}$ and $\mathcal{L}$, where $M\ll N$, namely, the eigenvectors corresponding to the largest eigenvalues in their absolute value. We order the eigenvalues in decreasing order from positive to negative. Let $\mathcal{PW}(M)$ be the space spanned by the $M$ leading eigenvectors of $\mathcal{L}$, called the Paley-Wiener space of $\mathcal{L}$. Let $p_{M}$ be the projection upon the Paley-Wiener space of $\mathcal{L}$. Let $R_{N}:L^{2}(\mathcal{S})\rightarrow L^{2}[N]$ be the sampling operator, defined by $(R_{N}f)_{j}=f(x_{j})$.
Let $f:\mathcal{S}\rightarrow{\mathbb{R}}$ be a bounded signal over the metric space and suppose that $f\in\mathcal{PW}(N)$. Denote $\mathbf{f}=(f(x_{n}))_{n=1}^{N}=R_{N}f$. By Monte Carlo theory, we have
[…]
So, in case the sampling is uniform. i.e. $w(x)=1$, we have
$L\approx N\mathcal{L},$
pointwise. To make this precise, let us recall Hoeffding’s Inequality.
Theorem A.1 (Hoeffding’s Inequality).
Let $Y_{1},\ldots,Y_{N}$ be independent random variables such that $a\leq Y_{n}\leq b$ almost surely. Then, for every $k>0$,
$\mathbb{P}\Big{(}\Big{|}\frac{1}{N}\sum_{n=1}^{N}(Y_{n}-\mathbb{E}[Y_{n}])\Big{|}\geq k\Big{)}\leq 2\exp\Big{(}-\frac{2k^{2}N}{(b-a)^{2}}\Big{)}.$
Using Hoeffding’s Inequality, one can now show that there is an event of probability more than $1-p$ in which for every $j\in[N]$, the error between $(\frac{1}{N}\mathbf{L}R_{N}f)_{j}$ and $\mathcal{L}f$ satisfies
$\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}R_{N}f\right\rVert={\mathcal{O}}\Big{(}M\sqrt{\frac{\log(1/p)+\log(N)+\log(2)}{N}}\Big{)}$
The fact that different graphs of different sizes $N$ approximate $\mathcal{L}$ with different scaling factors means that $\mathbf{L}$ is not value transferable. Let us show that $\mathbf{L}$ is index transferable.
We now evoke the transferability theorem – a slight reformulation of Section 3 of Theorem 4 in [Levie et al., 2021].
Theorem A.2.
Let $\mathcal{L}$ be a compact operator on $L^{2}(\mathcal{S})$ and $\mathbf{L}$ an operator on $\mathbb{C}^{N}$. Let $M\in\mathbb{N}$ and let $R_{M}$ be the sampling operator defined above. Let $\mathcal{PW}(M)$ be the space spanned by the $M$ leading eigenvectors of $\mathcal{L}$, and let $f\in\mathcal{PW}(M)$. Let $g:{\mathbb{R}}\rightarrow{\mathbb{R}}$ be Lipschitz with Lipschitz constant $D$. Then, there exists a constant $C$ (that depends on $M$) such that
$\left\lVert R_{N}g(\mathcal{L})f-\frac{1}{N}g({\mathbf{L}})R_{N}f\right\rVert\leq D\sqrt{M}\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}R_{N}f\right\rVert.$
For every leading eigenvalue $\lambda_{j}$ of $\mathcal{L}$ let $\gamma_{i_{j}}$ be the leading eigenvalue of $\mathcal{L}$ closest to $\lambda$, where, if there are repeated eigenvalues, $i_{j}$ is chosen arbitrarily out of the eigenvalues of $\mathbf{L}$ that best approximate $\lambda_{i}$.Let
$\delta=\min\{\alpha,\beta\}$
where
$\alpha=\min_{j\in[M]}\min\big{\{}\ \min_{i_{j}\neq i\in[N]}\left\lvert\lambda_{j}-\gamma_{i}\right\rvert\ ,\ \min_{j\neq m\in[M]}\left\lvert\lambda_{j}-\lambda_{m}\right\rvert\ ,\ \min_{j\neq m\in[M]}\left\lvert\gamma_{i_{j}}-\lambda_{m}\right\rvert\ \big{\}}$
and
$\beta=\min_{i\in[M],n\in[N]}\left\lvert\gamma_{i}-\gamma_{n}\right\rvert$
and suppose that $\delta>0$. For each $j\in[N]$ there exists a Lipschitz continuous function $g_{j}$ with Lipschitz constant $\delta^{-1}$ such that $g_{j}(\lambda_{j})=g_{j}(\gamma_{i_{j}})=1$ and $\gamma$ is zero on all other leading eigenvalues of $\mathcal{L}$ and all other eigenvalues of $\mathbf{L}$. Hence,
$g_{j}(\mathcal{L})f=p_{j}^{\mathcal{L}}f,\quad g_{j}(\mathbf{L})R_{N}f=p_{i_{j}}^{\mathbf{L}}\mathbf{f}$
where $p_{j}^{\mathcal{L}}$ is the projection upon the space spanned by the $j$-th eigenvector of $\mathcal{L}$, and $p_{i_{j}}^{\mathbf{L}}$ is the projection upon the space spanned by the $i_{j}$-th eigenvector of $\mathbf{L}$
Now, be the transferability theorem,
$\left\lVert R_{N}p^{\mathcal{L}}_{j}f-p^{\mathbf{L}}_{i_{j}}\mathbf{f}\right\rVert\leq D\sqrt{M}\left\lVert R_{N}\mathcal{L}f-\frac{1}{N}{\mathbf{L}}\mathbf{f}\right\rVert={\mathcal{O}}(\sqrt{M/N}).$
By induction over $j$, with base $j=1$, we must have $i_{j}=j$ for every $j\in[M]$.
Now, note that by standard Monte Carlo theory (evoking Hoeffding’s inequality again and intersecting events), we have
$\left\lvert\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}\right\rvert={\mathcal{O}}(M^{-1/2})$
in high probability. Hence, by the fact that
$\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}^{2}=\left(\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}\right)\left(\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}+\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}\right)$
and $\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}+\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}$ is bounded from below by the constant $\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}$,
$\left\lvert\left\lVert R_{N}p^{\mathcal{L}}_{j}f\right\rVert_{2}-\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}\right\rvert={\mathcal{O}}(M^{-1/2})$
This shows that
$\left\lVert p^{\mathcal{L}}_{j}f\right\rVert_{2}\approx\left\lVert p^{\mathbf{L}}_{j}\mathbf{f}\right\rVert_{2},$
which shows index transferability. Namely, for two graphs of $N$ and $N^{\prime}$ nodes sampled from $\mathcal{S}$, with corresponding Laplacians $\mathbf{L}$ and $\mathbf{L}^{\prime}$, by the triangle inequality,
$\left\lVert p^{\mathbf{L}}_{j}\mathbf{f}\right\rVert_{2}\approx\left\lVert p^{\mathbf{L}^{\prime}}_{j}\mathbf{f}\right\rVert_{2}.$
Next, we show value transferability for $\mathbf{N}$. Here,
$\mathbf{N}f(x_{j})\approx\frac{1}{V(x_{j})}\mathcal{L}f(x_{j}).$
Hence, if $V(x)=V$ is constant, we have $\mathbf{N}\approx\mathcal{L}$ up to a constant that does not depend on $N$. In this case, by a similar analysis to the above, $\mathbf{N}$ is value transferable. We note that $\mathbf{N}$ is also index transferable, but value transferability is guaranteed in a more general case, where we need not assume a separable spectrum.
Appendix B Additional Details on Experimental Study
We describe the experimental setups and additional details of our experiments in Sec. 5.
B.1 Semi-Supervised Node Classification
We provide a detailed overview of the experimental settings for semi-supervised node classification tasks, along with the validated hyperparameters used in our benchmarks.
Datasets. We consider six datasets in node classification tasks, including Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. The detailed statistics of the node classification benchmarks are summarized in Tab. LABEL:tab:node_classification_statistics. For datasets splitting on citation graphs (Cora, Citeseer, and Pubmed), we follow the standard splits from [Yang et al., 2016], using 20 nodes per class for training, 500 nodes for validation, and 1000 nodes for testing. For heterophilic graphs (Chameleon, Squirrel, and Actor), we adopt the sparse splitting method from [Chien et al., 2020, He et al., 2022], allocating 2.5% of samples for training, 2.5% for validation, and 95% for testing. The classification quality is assessed by computing the average classification accuracy across 10 random splits, along with a 95% confidence interval.
Baselines. For GCN, GAT, SAGE, ChebNet, ARMA, and APPNP, we use the implementations from the PyTorch Geometric library [Fey and Lenssen, 2019]. For other baselines, we use the implementations released by the respective authors.
Hyperparameters Settings. The hidden dimension is set to be either 64 or 128 for all models and datasets. We implement our proposed Node-level NLSFs using PyTorch and optimize the model with the Adam optimizer [Kingma and Ba, 2014]. To determine the optimal dropout probability, we search within the range $[0,0.9]$ in increments of $0.1$. The learning rate is examined within the set $\{1e^{-1},5e^{-2},1e^{-2},5e^{-3},1e^{-3}\}$. We explore weight decay values within the set $\{1e^{-2},5e^{-3},1e^{-3},5e^{-4},1e^{-4},5e^{-5},1e^{-5},5e^{-6},1e^{-6},0.0\}$. Furthermore, the number of layers is varied from 1 to 10. The number of leading eigenvectors $J$ in Index NLSFs is set within $[1,1e^{2}]$. The decay rate in the Value NLSFs is determined using dyadic sampling within the set $\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{3}{4}\}$, the sampling resolution $S$ within $[1,1e^{2}]$, and the number of the bands in Value NLSFs $K\leq S$. For hyperparameter optimization, we conduct a grid search using Optuna [Akiba et al., 2019] for each dataset. An early stopping criterion is employed during training, stopping the process if the validation loss does not decrease for 200 consecutive epochs.
Table 4: Node classification datasets statistics.
Dataset ##\## Nodes ##\## Classes ##\## Edges ##\## Features
Cora 2708 7 5278 1433
Citeseer 3327 6 4552 3703
Pubmed 19717 5 44324 500
Chameleon 2277 5 31371 2325
Squirrel 5201 5 198353 2089
Actor 7600 5 26659 932
B.2 Graph Classification
Table 5: Graph classification datasets statistics.
Dataset ##\## Graphs ($Z$) ##\## Classes ($C$) $\left\rvert N_{i}\right\rvert_{\text{min}}$ $\left\rvert N_{i}\right\rvert_{\text{max}}$ $\left\rvert N_{i}\right\rvert_{\text{avg}}$#\## Features ($d$)
MUTAG 188 2 10 28 17.93 7
PTC 344 2 2 64 14.29 18
ENZYMES 600 6 2 126 32.63 18
PROTEINS 1113 2 4 620 39.06 29
NCI1 4110 2 3 111 29.87 37
IMDB-B 1000 2 12 136 19.77 None
IMDB-M 1500 3 7 89 13.00 None
COLLAB 5000 3 32 492 74.50 None
We provide a comprehensive description of the experimental settings for graph classification tasks and the validated hyperparameters used in our benchmarks. The results are reported in the main paper.
Problem Setup. Consider a set of $Z$ graphs $\mathcal{G}_{Z}=\{G_{1},G_{2},\ldots,G_{Z}\}$, where in each graph $G_{i}=([N_{i}],\mathcal{E}_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, we have $N_{i}$ nodes for each graph $G_{i}$, $\mathcal{E}_{i}$ represents the edge set, $\mathbf{A}_{i}\in\mathbb{R}^{N_{i}\times N_{i}}$ denotes the edge weights, and $\mathbf{X}_{i}\in\mathbb{R}^{N_{i}\times d}$ represents the node feature matrix with $d$-dimensional node attributes. Let $\mathbf{Y}_{Z}\in\mathbb{R}^{Z\times C}$ be the label matrix with $C$ classes such that $y_{i,j}=1$ if the graph $G_{i}$ belongs to the class $j$, and $y_{i,j}=0$ otherwise. Given a set of $Z^{\prime}$ graphs $\mathcal{G}_{Z^{\prime}}\subset\mathcal{G}$, where $Z^{\prime}<Z$, with the label information $\mathbf{Y}_{Z^{\prime}}\in\mathbb{R}^{Z^{\prime}\times C}$, our goal is to classify the set of unseen graph labels of $\mathcal{G}_{Z}\setminus\mathcal{G}_{Z^{\prime}}$.
Datasets. We consider eight datasets [Morris et al., 2020] for graph classification tasks, including five bioinformatics: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, and three social network datasets: IMDB-B, IMDB-M, and COLLAB. The detailed statistics of the graph classification benchmarks are summarized in Tab. LABEL:tab:graph_classification_statistics. We use the random split from [Velickovic et al., 2019, Ying et al., 2018, Ma et al., 2019, Zhu et al., 2021b], using 80% for training, 10% for validation, and 10% for testing. This process is repeated 10 times, and we report the average performance and standard deviation.
Baselines. For GCN, GAT, SAGE, ChebNet, ARMA, APPNP, and DiffPool, we use the implementations from the PyTorch Geometric library [Fey and Lenssen, 2019]. For other baselines, we use the implementations released by the respective authors.
Hyperparameters Settings. The dimension of node representations is set to 128 for all methods and datasets. We implement the proposed Pooling-NLSFs and Graph-level NLSFs using PyTorch and optimize the model with the Adam optimizer [Kingma and Ba, 2014]. A readout function is applied to aggregate the node representations for each graph, utilizing mean, add, max, or RMS poolings. The learning rate and weight decay are searched within $\{1e^{-1},1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$, the pooling ratio within $[0.1,0.9]$ with step $0.1$, the number of layers within $[1,10]$ with step $1$, the number of leading eigenvectors $J$ in Index NLSFs within $[1,1e^{2}]$, the decay rate in the Value NLSFs using dyadic sampling within $\{\frac{1}{4},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{3}{4}\}$, the sampling resolution $S$ within $[1,1e^{2}]$, and the number of the bands in Value NLSFs within $K\leq S$. The graph-level representation is then fed into an MLP with a (log)softmax classifier, using a cross-entropy loss function for predictions over the labels. Specifically, the MLP consists of three fully connected layers with 256, 128, and 64 neurons, respectively, followed by a (log)softmax classifier. We conduct a grid search on the hyperparameters for each dataset using Optuna [Akiba et al., 2019]. An early stopping criterion is employed during training, stopping the process if the validation loss does not decrease for 100 consecutive epochs.
Appendix C Additional Experimental Results
Here, we present additional experiments on node and graph classification benchmarks, ablation studies, runtime analysis, and uniform sub-bands.
C.1 Semi-Supervised Node Classification Following [He et al., 2022] Protocol
Table 6: Semi-supervised node classification accuracy with random split, following the experimental protocol established by [He et al., 2022]. Results marked with are taken from [He et al., 2022].
Cora Citeseer Pubmed Chameleon Squirrel Actor
GCN 79.19$\pm$1.4 69.71$\pm$1.3 78.81$\pm$0.8 38.15$\pm$3.8 31.18$\pm$1.0 22.74$\pm$2.3
GAT 80.03$\pm$0.8 68.16$\pm$0.9 77.26$\pm$0.5 34.16$\pm$1.2 27.40$\pm$1.4 24.35$\pm$1.7
SAGE 72.68$\pm$1.9 63.87$\pm$1.2 77.68$\pm$0.8 31.77$\pm$1.8 22.67$\pm$1.8 25.61$\pm$1.8
ChebNet 78.08$\pm$0.9 67.87$\pm$1.5 73.96$\pm$1.7 37.15$\pm$1.5 26.55$\pm$0.5 26.58$\pm$1.9
ChebNetII 82.42$\pm$0.6 69.89$\pm$1.2 79.53$\pm$1.0 43.42$\pm$3.5 33.96$\pm$1.2 30.18$\pm$0.8
CayleyNet 80.25$\pm$1.4 66.46$\pm$2.9 75.42$\pm$3.4 34.52$\pm$3.1 24.08$\pm$2.9 27.42$\pm$3.3
APPNP 82.39$\pm$0.7 69.79$\pm$0.9 79.97$\pm$1.6 32.73$\pm$2.3 24.50$\pm$0.9 29.74$\pm$1.0
GPRGNN 82.37$\pm$0.9 69.22$\pm$1.3 79.28$\pm$1.3 33.03$\pm$1.9 24.36$\pm$1.5 28.58$\pm$1.0
ARMA 79.14$\pm$1.1 69.35$\pm$1.4 78.31$\pm$1.3 37.42$\pm$1.7 24.15$\pm$0.9 27.02$\pm$2.3
$\mathtt{att}$-Node-level NLSFs 82.94$\pm$1.1 72.13$\pm$1.1 79.62$\pm$1.2 46.75$\pm$1.3 35.17$\pm$1.6 29.96$\pm$0.8
We present additional experimental results for semi-supervised node classification using random splits, adhering to the protocol established by [He et al., 2022]. The results, summarized in Tab. LABEL:tab:node_classification_accuracy_following_chebnetII_setup, demonstrate the classification accuracy across six benchmark datasets: Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. Our $\mathtt{att}$-Node-level NLSFs achieve the highest accuracy on four out of the six datasets, outperforming other models significantly. On Pubmed, it records a close second-highest accuracy, slightly behind APPNP. Our method achieves a competitive second place in the Actor dataset. $\mathtt{att}$-Node-level NLSFs demonstrate substantial improvements, particularly in challenging datasets like Chameleon and Squirrel. The comparison models, including GCN, GAT, SAGE, ChebNet, ChebNetII, CayleyNet, APPNP, GPRGNN, and ARMA, varied in their effectiveness, with ChebNetII and APPNP also showing strong results on several datasets. These findings highlight the efficacy of the $\mathtt{att}$-Node-level NLSFs in semi-supervised node classification tasks.
C.2 Node Classification on Filtered Chameleon and Squirrel Datasets in Dense Split Setting
Table 7: Full Performance comparison of node classification on original and filtered Chameleon and Squirrel datasets in dense split setting. The baseline results used for comparison are taken from [Platonov et al., 2023].
Chameleon Squirrel
Original Filtered Original Filtered
ResNet 49.52$\pm$1.7 36.73$\pm$4.7 33.88$\pm$1.8 36.55$\pm$1.8
ResNet+SGC 49.93$\pm$2.3 41.01$\pm$4.5 34.36$\pm$1.2 38.36$\pm$2.0
ResNet+adj 71.07$\pm$2.2 38.67$\pm$3.9 65.46$\pm$1.6 38.37$\pm$2.0
GCN 50.18$\pm$3.3 40.89$\pm$4.1 39.06$\pm$1.5 39.47$\pm$1.5
SAGE 50.18$\pm$1.8 37.77$\pm$4.1 35.83$\pm$1.3 36.09$\pm$2.0
GAT 45.02$\pm$1.8 39.21$\pm$3.1 32.21$\pm$1.6 35.62$\pm$2.1
GAT-sep 50.24$\pm$2.2 39.26$\pm$2.5 35.72$\pm$2.0 35.46$\pm$3.1
GT 44.93$\pm$1.4 38.87$\pm$3.7 31.61$\pm$1.1 36.30$\pm$2.0
GT-sep 50.33$\pm$2.6 40.31$\pm$3.0 36.08$\pm$1.6 36.66$\pm$1.6
H2GCN 46.27$\pm$2.7 26.75$\pm$3.6 29.45$\pm$1.7 35.10$\pm$1.2
CPGNN 48.77$\pm$2.1 33.00$\pm$3.2 30.91$\pm$2.0 30.04$\pm$2.0
GPRGNN 47.26$\pm$1.7 39.93$\pm$3.3 33.39$\pm$2.1 38.95$\pm$2.0
FSGNN 77.85$\pm$0.5 40.61$\pm$3.0 68.93$\pm$1.7 35.92$\pm$1.3
GloGNN 70.04$\pm$2.1 25.90$\pm$3.6 61.21$\pm$2.0 35.11$\pm$1.2
FAGCN 64.23$\pm$2.0 41.90$\pm$2.7 47.63$\pm$1.9 41.08$\pm$2.3
GBKGNN 51.36$\pm$1.8 39.61$\pm$2.6 37.06$\pm$1.2 35.51$\pm$1.7
JacobiConv 68.33$\pm$1.4 39.00$\pm$4.2 46.17$\pm$4.3 29.71$\pm$1.7
$\mathtt{att}$-Node-level NLSFs 79.42$\pm$1.6 42.06$\pm$1.3 67.81$\pm$1.4 42.18$\pm$1.2
Following [Platonov et al., 2023], we conduct additional experiments on the original and filtered Chameleon and Squirrel datasets in the dense split setting. We use the same random splits as in [Platonov et al., 2023], dividing the datasets into 48% for training, 32% for validation, and 20% for testing. We compare the Node-level NLSFs using Laplacian attention with GCN [Kipf and Welling, 2016], SAGE [Hamilton et al., 2017], GAT [Veličković et al., 2017], GT [Shi et al., 2020], H2GCN [Zhu et al., 2020], CPGNN [Zhu et al., 2021a], GPRGNN [Chien et al., 2020], FSGNN [Maurya et al., 2022], GloGNN [Li et al., 2022], FAGCN [Bo et al., 2021], GBKGNN [Du et al., 2022], JacobiConv [Wang and Zhang, 2022], and ResNet [He et al., 2016] with GNN models [Platonov et al., 2023]. The study by [Zhu et al., 2020] demonstrates the benefits of separating ego- and neighbor-embeddings in the GNN aggregation step when dealing with heterophily. Therefore, [Platonov et al., 2023] also adopts this approach for the GNN aggregation step in GAT and GT models, denoted as “sep.” The baseline results used for comparison are taken from [Platonov et al., 2023]. Tab. 7 presents the full performance comparison on the original and filtered datasets. Note that Tab. 7 is the same as Tab. 2 but with more baseline methods. $\mathtt{att}$-Node-level NLSFs achieve the highest accuracy on both the filtered Chameleon and Squirrel datasets. Additionally, $\mathtt{att}$-Node-level NLSFs demonstrate strong performance on the original Chameleon dataset, achieving the highest accuracy and the second-highest accuracy on the original Squirrel dataset. $\mathtt{att}$-Node-level NLSFs show less sensitivity to node duplicates and exhibit stronger generalization ability, further validating the reliability of the Chameleon and Squirrel datasets in the dense split setting.
C.3 Ablation Study on $\mathtt{att}$-Node-level NLSFs, $\mathtt{att}$-Graph-NLSF, and $\mathtt{att}$-Pooling-NLSFs
Table 8: An ablation study investigated the effect of Node-level NLSFs on node classification, comparing the use of Index NLSFs and Value NLSFs. The symbol ()($\uparrow$)( ↑ ) denotes an improvement using Laplacian attention.
Cora Citeseer Pubmed Chameleon Squirrel Actor
$\Theta_{\text{ind}}(\mathbf{L},\cdot)$ 82.46$\pm$1.2 71.31$\pm$1.0 80.97$\pm$0.9 44.37$\pm$1.8 34.12$\pm$1.4 29.88$\pm$1.1
$\Theta_{\text{ind}}(\mathbf{N},\cdot)$ 81.73$\pm$1.4 70.25$\pm$0.8 79.88$\pm$1.2 44.52$\pm$1.7 34.26$\pm$1.5 29.97$\pm$1.7
$\Theta_{\text{val}}(\mathbf{L},\cdot)$ 80.25$\pm$1.5 70.43$\pm$1.7 80.06$\pm$1.6 45.52$\pm$1.2 35.23$\pm$0.8 32.69$\pm$1.9
$\Theta_{\text{val}}(\mathbf{N},\cdot)$ 81.98$\pm$0.7 71.16$\pm$1.3 80.33$\pm$1.7 47.91$\pm$1.4 35.78$\pm$0.8 33.53$\pm$1.4
$\mathtt{att}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{L},\cdot))$ 82.46$\pm$1.2 71.48$\pm$0.8 ()($\uparrow$)( ↑ ) 80.97$\pm$0.9 46.71$\pm$2.2 ()($\uparrow$)( ↑ ) 36.92$\pm$1.1 ()($\uparrow$)( ↑ ) 32.69$\pm$1.9
$\mathtt{att}(\Theta_{\text{ind}}(\mathbf{N},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ 81.98$\pm$0.7 72.45$\pm$1.4 ()($\uparrow$)( ↑ ) 80.33$\pm$1.7 47.91$\pm$1.4 36.87$\pm$0.7 ()($\uparrow$)( ↑ ) 33.53$\pm$1.4
$\mathtt{att}(\Theta_{\text{ind}}(\mathbf{N},\cdot),\Theta_{\text{val}}(\mathbf{L},\cdot))$ 82.65$\pm$1.2 ()($\uparrow$)( ↑ ) 71.26$\pm$1.7 ()($\uparrow$)( ↑ ) 80.56$\pm$1.7 ()($\uparrow$)( ↑ ) 45.98$\pm$2.3 ()($\uparrow$)( ↑ ) 37.03$\pm$1.9 ()($\uparrow$)( ↑ ) 33.09$\pm$1.2 ()($\uparrow$)( ↑ )
$\mathtt{att}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ 84.75$\pm$0.7 ()($\uparrow$)( ↑ ) 73.62$\pm$1.1 ()($\uparrow$)( ↑ ) 81.93$\pm$1.0 ()($\uparrow$)( ↑ ) 49.68$\pm$1.6 ()($\uparrow$)( ↑ ) 38.25$\pm$0.7 ()($\uparrow$)( ↑ ) 34.72$\pm$0.9 ()($\uparrow$)( ↑ )
Table 9: An ablation study investigated the effect of Graph-NLSFs on graph classification, comparing the use of Index NLSFs and Value NLSFs. The symbol ()($\uparrow$)( ↑ ) denotes an improvement using Laplacian attention.
MUTAG PTC ENZYMES PROTEINS NCI1 IMDB-B IMDB-M COLLAB
$\Phi_{\text{ind}}(\mathbf{L},\cdot)$ 82.14$\pm$1.2 65.73$\pm$2.4 62.31$\pm$1.4 78.22$\pm$1.6 80.51$\pm$1.2 63.27$\pm$1.1 50.29$\pm$1.4 70.06$\pm$1.3
$\Phi_{\text{ind}}(\mathbf{N},\cdot)$ 83.27$\pm$0.3 66.08$\pm$2.1 60.18$\pm$0.5 80.31$\pm$1.4 78.40$\pm$0.6 64.58$\pm$2.2 52.33$\pm$0.9 71.88$\pm$1.7
$\Phi_{\text{val}}(\mathbf{L},\cdot)$ 83.40$\pm$1.4 65.67$\pm$1.4 66.42$\pm$1.1 83.36$\pm$1.4 76.17$\pm$1.2 72.71$\pm$1.4 52.13$\pm$0.9 74.19$\pm$1.0
$\Phi_{\text{val}}(\mathbf{N},\cdot)$ 81.19$\pm$0.3 66.20$\pm$0.7 63.32$\pm$1.2 78.20$\pm$0.9 78.68$\pm$1.2 74.26$\pm$1.8 52.49$\pm$0.7 79.06$\pm$1.2
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{L},\cdot),\Phi_{\text{val}}(\mathbf{L},\cdot))$ 83.40$\pm$1.4 66.32$\pm$1.9 ()($\uparrow$)( ↑ ) 66.42$\pm$1.1 83.36$\pm$1.4 80.51$\pm$1.2 72.71$\pm$1.4 52.13$\pm$0.9 75.14$\pm$1.8 ()($\uparrow$)( ↑ )
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{N},\cdot),\Phi_{\text{val}}(\mathbf{N},\cdot))$ 83.71$\pm$1.7 ()($\uparrow$)( ↑ ) 67.14$\pm$1.5 ()($\uparrow$)( ↑ ) 65.97$\pm$1.1 ()($\uparrow$)( ↑ ) 81.79$\pm$1.2 ()($\uparrow$)( ↑ ) 79.83$\pm$0.6 ()($\uparrow$)( ↑ ) 74.26$\pm$1.8 52.49$\pm$0.7 79.06$\pm$1.2
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{N},\cdot),\Phi_{\text{val}}(\mathbf{L},\cdot))$ 83.78$\pm$0.8 ()($\uparrow$)( ↑ ) 66.20$\pm$0.7 66.42$\pm$1.1 83.36$\pm$1.4 78.40$\pm$0.6 72.71$\pm$1.4 52.33$\pm$0.9 76.22$\pm$0.8 ()($\uparrow$)( ↑ )
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{L},\cdot),\Phi_{\text{val}}(\mathbf{N},\cdot))$ 84.13$\pm$1.5 ()($\uparrow$)( ↑ ) 68.17$\pm$1.0 ()($\uparrow$)( ↑ ) 65.94$\pm$1.6 ()($\uparrow$)( ↑ ) 82.69$\pm$1.9 ()($\uparrow$)( ↑ ) 80.51$\pm$1.2 74.26$\pm$1.8 52.49$\pm$0.7 79.06$\pm$1.2
Table 10: An ablation study investigated the effect of Pooling-NLSFs on graph classification, comparing the use of Index NLSFs and Value NLSFs. The symbol ()($\uparrow$)( ↑ ) denotes an improvement using Laplacian attention.
MUTAG PTC ENZYMES PROTEINS NCI1 IMDB-B IMDB-M COLLAB
$\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$ 86.41$\pm$1.9 68.76$\pm$0.9 69.88$\pm$1.3 84.27$\pm$1.3 80.33$\pm$0.8 60.40$\pm$1.3 51.01$\pm$1.3 71.28$\pm$0.9
$\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$ 84.52$\pm$0.8 67.19$\pm$1.2 66.37$\pm$2.1 81.12$\pm$1.7 77.05$\pm$2.2 62.08$\pm$1.6 52.48$\pm$1.0 72.03$\pm$1.2
$\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$ 83.64$\pm$0.9 67.74$\pm$1.3 65.87$\pm$1.4 84.18$\pm$1.4 79.43$\pm$1.4 71.98$\pm$2.0 51.85$\pm$1.7 78.80$\pm$1.1
$\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$ 86.58$\pm$1.0 69.13$\pm$1.1 68.89$\pm$0.4 84.80$\pm$1.0 80.82$\pm$0.8 76.48$\pm$1.8 54.12$\pm$1.0 81.31$\pm$0.7
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{L},\cdot))$ 86.41$\pm$1.9 68.76$\pm$0.9 69.88$\pm$1.3 84.27$\pm$1.3 80.33$\pm$0.8 72.44$\pm$1.3 ()($\uparrow$)( ↑ ) 51.85$\pm$1.7 78.80$\pm$1.1
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{N},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ 86.58$\pm$1.0 69.67$\pm$1.2 ()($\uparrow$)( ↑ ) 69.06$\pm$1.1 ()($\uparrow$)( ↑ ) 84.80$\pm$1.0 80.82$\pm$0.8 76.48$\pm$1.8 54.37$\pm$1.2 ()($\uparrow$)( ↑ ) 81.70$\pm$0.9 ()($\uparrow$)( ↑ )
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{N},\cdot),\Theta_{\text{val}}(\mathbf{L},\cdot))$ 84.93$\pm$1.7 ()($\uparrow$)( ↑ ) 68.14$\pm$1.6 ()($\uparrow$)( ↑ ) 67.26$\pm$1.0 ()($\uparrow$)( ↑ ) 84.18$\pm$1.4 79.43$\pm$1.4 72.64$\pm$1.2 ()($\uparrow$)( ↑ ) 52.48$\pm$1.0 78.80$\pm$1.1
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ 86.89$\pm$1.2 ()($\uparrow$)( ↑ ) 71.02$\pm$1.3 ()($\uparrow$)( ↑ ) 69.94$\pm$1.0 ()($\uparrow$)( ↑ ) 84.89$\pm$0.9 ()($\uparrow$)( ↑ ) 80.95$\pm$1.4 ()($\uparrow$)( ↑ ) 76.78$\pm$1.9 ()($\uparrow$)( ↑ ) 55.28$\pm$1.7 ()($\uparrow$)( ↑ ) 82.19$\pm$1.3 ()($\uparrow$)( ↑ )
In Sec. 3.5, we note that using the graph Laplacian $\mathbf{L}$ in Index NLSFs and the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs is transferable. Since real-world graphs often fall between these two boundary cases, we present the Laplacian attention NLSFs that operate between them at both the node-level and graph-level. Indeed, as demonstrated in Sec. 5, the proposed $\mathtt{att}$-Node-level NLSFs, $\mathtt{att}$-Graph-level NLSFs, and $\mathtt{att}$-Pooling-NLSFs outperform existing spectral GNNs.
We conduct an ablation study to evaluate the contribution and effectiveness of different components within the $\mathtt{att}$-Node-level NLSFs, $\mathtt{att}$-Graph-level-NLSFs, and $\mathtt{att}$-Pooling-NLSFs on node and graph classification tasks. Specifically, we compare the Index NLSFs and Value NLSFs using both the graph Laplacian $\mathbf{L}$ and the normalized graph Laplacian $\mathbf{N}$ to understand their individual and Laplacian attention impact on these tasks.
The ablation study of $\mathtt{att}$-Node-level NLSPs for node classification is summarized in Tab. LABEL:tab:node_classification_nlsfp_ablation. We investigate the performance on six node classification benchmarks as in Sec. 5, including Cora, Citeseer, Pubmed, Chameleon, Squirrel, and Actor. Tab. LABEL:tab:node_classification_nlsfp_ablation shows that using the graph Laplacian $\mathbf{L}$ in Index NLSFs (denoted as $\Theta_{\text{ind}}(\mathbf{L},\cdot)$) and the normalized graph Laplacian $\mathbf{N}$ in Value NLSFs (denoted as $\Theta_{\text{val}}(\mathbf{N},\cdot)$ ) has superior node classification accuracy compared to using the normalized graph Laplacian $\mathbf{N}$ in Index NLSFs (denoted as $\Theta_{\text{ind}}(\mathbf{N},\cdot)$ ) and the graph Laplacian $\mathbf{L}$ in Value NLSFs (denoted as $\Theta_{\text{val}}(\mathbf{L},\cdot)$). This is in line with our theoretical findings in App. A.5. Moreover, the $\mathtt{att}$-Node-level NLSFs using the Laplacian attention $\mathtt{att}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ yield the highest accuracies across all datasets, corroborating the findings in Sec. 3.5. We also note that without Laplacian attention, the Node-level NLSFs ($\Theta_{\text{ind}}(\mathbf{L},\cdot)$ and $\Theta_{\text{val}}(\mathbf{N},\cdot)$) alone still achieve more effective classification performance compared to existing baselines, as shown in Tab. LABEL:tab:node_classification_accuracy.
Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf demonstrates the ablation study of Graph-level NLSFs for graph classification tasks. We examine the eight graph datasets as in Sec. 5, including MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB. Similar to the above, we investigate the Index and Value settings using graph Laplacian $\mathbf{L}$ and normalized graph Laplacian $\mathbf{N}$, including $\Phi_{\text{ind}}(\mathbf{L},\cdot)$, $\Phi_{\text{ind}}(\mathbf{N},\cdot)$, $\Phi_{\text{val}}(\mathbf{L},\cdot)$, and $\Phi_{\text{val}}(\mathbf{N},\cdot)$, along with their variants using Laplacian attention. Here, we see that $\mathtt{att}$-Graph-level NLSFs do not show significant improvement over the standard Graph-level NLSFs. Notably, $\Phi_{\text{val}}(\mathbf{N},\cdot)$ outperforms other models in social network datasets (IMDB-B, IMDB-M, and COLLAB), where the node features are augmented by the node degree. We plan to investigate the limited graph attribution in future work.
The ablation study of $\mathtt{att}$-Pooling-NLSFs for graph classification is reported in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling. Similar to Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf, we consider eight graph classification benchmarks: MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB. Tabl. LABEL:tab:graph_classification_nlsfp_ablation_pooling demonstrates that using the graph Laplacian $\mathbf{L}$ in Index Pooling-NLSFs (denoted as $\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$) and the normalized graph Laplacian $\mathbf{N}$ in Value Pooling-NLSFs (denoted as $\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$) achieves superior graph classification accuracy compared to using the normalized graph Laplacian $\mathbf{N}$ in Index Pooling-NLSFs (denoted as $\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$) and the graph Laplacian $\mathbf{L}$ in Value Pooling-NLSFs (denoted as $\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$). This finding aligns with our theoretical results in App. A.5. Moreover, the $\mathtt{att}$-Pooling-NLSFs using the Laplacian attention $\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ yield the highest accuracies across all datasets, corroborating the findings in Sec. 3.5. In addition, $\mathtt{att}$-Pooling-NLSFs consistently outperform $\mathtt{att}$-Graph-level NLSFs as shown in Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf, indicating that the node features learned in our Node-level NLSFs representation are more expressive. This supports our theoretical findings in Sec. 4.3.
The ablation study demonstrates that the Laplacian attention between the Index and Value NLSFs significantly enhances classification accuracy for both node and graph classification tasks across various datasets, outperforming existing baselines.
C.4 Runtime Analysis
Table 11: Average running time per epoch(ms)/average total running time(s).
Cora Citeseer Pubmed Chameleon Squirrel Actor
GCN 8.97/2.1 9.1/2.3 12.15/3.9 11.68/2.7 25.52/6.2 14.51/3.8
GAT 13.13/3.3 13.87/3.6 19.42/6.2 14.83/3.4 46.13/15.9 20.13/4.4
SAGE 11.72/2.2 12.11/2.4 25.31/6.1 60.61/12.7 321.65/72.8 25.16/5.4
ChebNet 21.36/4.9 22.51/5.3 34.53/13.1 42.21/16.0 38.21/45.1 42.91/9.3
ChebNetII 20.53/5.9 20.61/5.7 33.57/12.9 39.03/17.3 37.29/38.04 40.05/9.1
CayleyNet 401.24/79.3 421.69/83.4 723.61/252.3 848.51/389.4 972.53/361.8 794.61/289.4
APPNP 18.31/4.2 19.17/4.8 19.63/5.9 18.56/3.8 24.18/4.9 15.93/4.6
GPRGNN 19.07/3.8 18.69/4.0 19.77/6.3 19.31/3.6 28.31/5.5 17.28/4.8
ARMA 20.91/5.2 19.33/4.9 34.27/14.5 41.63/19.7 39.42/42.7 46.22/5.7
$\mathtt{att}$-Node-level NLSFs 18.22/4.5 18.51/4.4 20.23/6.1 28.51/17.1 25.56/5.1 17.09/4.6
In our NLSFs, the eigendecomposition can be calculated once for each graph and reused in the training process. This step is essential as the cost of the forward pass during model training often surpasses the initial eigendecomposition preprocessing cost. Note that the computation time for eigendecomposition is considerably less than the time needed for model training. For medium and large graphs, the computation time is further reduced when partial eigendecomposition is utilized, making it more efficient than the training times of competing baselines. To evaluate the computational complexity of our NLSFs compared to baseline models, we report the empirical training time in Tab. LABEL:tab:runtime. Our Node-level NLSFs showcase competitive running times, with moderate values per epoch and total running times that are comparable to the most efficient models like GCN, GPRGNN, and APPNP. Notably, Node-level NLSFs are particularly efficient for the Cora, Citeseer, and Pubmed datasets. For the dense Squirrel graph, our Node-level NLSFs exhibit efficient performance with a moderate running time, outperforming several models that struggle with significantly higher times.
C.5 Diagonal Nonlinear Spectral Filter
Table 12: Graph classification performance using Diagonal NLSFs, including index-by-index Index-NLSFs and band-by-band Value-NLSFs, along with their variants using Laplacian attention. The symbol ()($\uparrow$)( ↑ ) denotes an improvement using Laplacian attention.
MUTAG PTC ENZYMES PROTEINS NCI1 IMDB-B IMDB-M COLLAB
$\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$ 82.33$\pm$1.7 66.43$\pm$1.8 69.69$\pm$2.4 83.47$\pm$2.2 77.43$\pm$1.4 60.17$\pm$0.8 50.04$\pm$1.2 70.89$\pm$0.9
$\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$ 82.85$\pm$0.8 63.40$\pm$1.7 68.06$\pm$2.1 82.69$\pm$0.8 78.00$\pm$1.4 62.27$\pm$1.8 51.74$\pm$1.4 70.69$\pm$1.3
$\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$ 82.77$\pm$0.8 63.20$\pm$0.9 64.19$\pm$2.3 83.39$\pm$1.7 78.26$\pm$1.4 70.61$\pm$2.2 51.86$\pm$1.9 77.03$\pm$1.5
$\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$ 80.29$\pm$0.8 65.02$\pm$1.4 65.29$\pm$1.7 81.62$\pm$1.2 78.28$\pm$1.0 74.33$\pm$1.7 52.89$\pm$0.9 80.41$\pm$0.8
$\mathtt{att}^{\mathtt{P}}(\Gamma_{\text{ind}}(\mathbf{L},\cdot),\Gamma_{\text{val}}(\mathbf{L},\cdot))$ 83.19$\pm$1.4 ()($\uparrow$)( ↑ ) 66.43$\pm$1.8 69.69$\pm$2.4 83.47$\pm$2.2 78.26$\pm$1.4 70.61$\pm$2.2 52.03$\pm$1.4 ()($\uparrow$)( ↑ ) 77.56$\pm$1.3 ()($\uparrow$)( ↑ )
$\mathtt{att}^{\mathtt{P}}(\Gamma_{\text{ind}}(\mathbf{N},\cdot),\Gamma_{\text{val}}(\mathbf{N},\cdot))$ 83.34$\pm$1.2 ()($\uparrow$)( ↑ ) 66.58$\pm$1.2 ()($\uparrow$)( ↑ ) 68.26$\pm$1.9 ()($\uparrow$)( ↑ ) 82.99$\pm$1.8 ()($\uparrow$)( ↑ ) 78.28$\pm$1.0 74.33$\pm$1.7 53.14$\pm$0.8 ()($\uparrow$)( ↑ ) 80.41$\pm$0.8
$\mathtt{att}^{\mathtt{P}}(\Gamma_{\text{ind}}(\mathbf{N},\cdot),\Gamma_{\text{val}}(\mathbf{L},\cdot))$ 83.42$\pm$1.5 ()($\uparrow$)( ↑ ) 64.08$\pm$1.7 ()($\uparrow$)( ↑ ) 68.06$\pm$2.1 83.39$\pm$1.7 78.26$\pm$1.4 70.61$\pm$2.2 52.13$\pm$1.4 ()($\uparrow$)( ↑ ) 78.12$\pm$1.9 ()($\uparrow$)( ↑ )
$\mathtt{att}^{\mathtt{P}}(\Gamma_{\text{ind}}(\mathbf{L},\cdot),\Gamma_{\text{val}}(\mathbf{N},\cdot))$ 83.85$\pm$1.4 ()($\uparrow$)( ↑ ) 67.12$\pm$1.6 ()($\uparrow$)( ↑ ) 69.69$\pm$2.4 83.47$\pm$2.2 78.28$\pm$1.0 74.33$\pm$1.7 53.91$\pm$1.6 ()($\uparrow$)( ↑ ) 81.03$\pm$1.2 ()($\uparrow$)( ↑ )
Our primary objective in this work is to introduce new GNNs that are equivariant to functional symmetries, based on a novel spectral domain that is transferable between graphs. We emphasize the unique aspects of our method rather than providing an exhaustive list of operators in this group, which, while important, is a key direction for future research.
Here, we present a new type of NLSFs, termed Diagonal NLSFs, which are index-by-index Index-NLSFs and band-by-band Value-NLSFs. We denote them as follows:
$\Gamma_{\text{ind}}({\bm{\Delta}},\mathbf{X})=\sum_{j=1}^{J}\gamma_{j}\left(\left\lVert\mathbf{P}_{j}\mathbf{X}\right\rVert_{\rm sig}\right)\frac{\mathbf{P}_{j}\mathbf{X}}{\left\lVert\mathbf{P}_{j}\mathbf{X}\right\rVert^{a}_{\rm sig}+e}\text{ and }$
$\Gamma_{\text{val}}({\bm{\Delta}},\mathbf{X})=\sum_{j=1}^{K}\gamma_{j}\left(\left\lVert g_{j}({\bm{\Delta}})\mathbf{X}\right\rVert_{\rm sig}\right)\frac{g_{j}(\mathbf{L})\mathbf{X}}{\left\lVert g_{p}({\bm{\Delta}})\mathbf{X}\right\rVert^{a}_{\rm sig}+e},$
where $a,e$ are as before. Note that these operators are also equivariant to our group actions.
We investigate Diagonal NLSFs in graph classification tasks as described in Sec. 5, including MUTAG, PTC, ENZYMES, PROTEINS, NCI1, IMDB-B, IMDB-M, and COLLAB datasets. Unlike Graph-NLSFs, which are fully spectral and map a sequence of frequency coefficients to an output vector, Diagonal NLSFs do not possess such a sequential spectral form. In Diagonal NLSFs, the index-by-index (or band-by-band) frequency response is projected back to the graph domain. Consequently, for graph classification tasks, we apply the readout function as defined for Pooling-NLSFs in Sec. 3.4.
Following App. C.3, we examine index-by-index Index and band-by-band Value settings using graph Laplacian $\mathbf{L}$ and normalized graph Laplacian $\mathbf{N}$, including $\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$, $\Gamma_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$, $\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$, and $\Gamma_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$, along with their variants using Laplacian attention, where $\mathtt{P}$ denotes the pooling function as in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling.
Tab. LABEL:tab:graph_classification_nlsfp_diagonal_nlsf presents the graph classification accuracy using Diagonal NLSFs. It shows that incorporating Laplacian attention consistently improves classification performance, aligning with the findings in App. C.3. We note that Diagonal-NLSFs perform comparably to existing baselines in graph classification benchmarks. However, Diagonal-NLSFs are generally less effective compared to Pooling-NLSFs, as shown in Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling.
C.6 Uniform Sub-Bands
Figure 3: Illustration of dyadic sub-bands for $r=\frac{1}{2}$ and $S=4$.
Table 13: Graph classification performance using Graph-NLSFs with uniform sub-bands. The symbol ()($\uparrow$)( ↑ ) denotes an improvement using Laplacian attention.
MUTAG PTC ENZYMES PROTEINS NCI1 IMDB-B IMDB-M COLLAB
$\Phi_{\text{ind}}(\mathbf{L},\cdot)$ 82.14$\pm$1.2 65.73$\pm$2.4 62.31$\pm$1.4 78.22$\pm$1.6 80.51$\pm$1.2 63.27$\pm$1.1 50.29$\pm$1.4 70.06$\pm$1.3
$\Phi_{\text{ind}}(\mathbf{N},\cdot)$ 83.27$\pm$0.3 66.08$\pm$2.1 60.18$\pm$0.5 80.31$\pm$1.4 78.40$\pm$0.6 64.58$\pm$2.2 52.33$\pm$0.9 71.88$\pm$1.7
$\Phi_{\text{val}}(\mathbf{L},\cdot)$ 81.73$\pm$1.4 64.81$\pm$1.4 64.46$\pm$1.2 78.31$\pm$0.9 78.22$\pm$1.4 70.84$\pm$2.3 50.42$\pm$1.8 73.49$\pm$1.4
$\Phi_{\text{val}}(\mathbf{N},\cdot)$ 81.12$\pm$0.6 65.13$\pm$0.7 60.06$\pm$1.3 78.88$\pm$0.4 78.48$\pm$1.3 74.05$\pm$2.4 51.13$\pm$0.8 78.16$\pm$1.3
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{L},\cdot),\Phi_{\text{val}}(\mathbf{L},\cdot))$ 82.76$\pm$0.9 ()($\uparrow$)( ↑ ) 65.73$\pm$2.4 64.58$\pm$1.7 ()($\uparrow$)( ↑ ) 79.23$\pm$1.2 ()($\uparrow$)( ↑ ) 80.51$\pm$1.2 70.84$\pm$2.3 51.04$\pm$2.2 ()($\uparrow$)( ↑ ) 74.21$\pm$1.2 ()($\uparrow$)( ↑ )
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{N},\cdot),\Phi_{\text{val}}(\mathbf{N},\cdot))$ 83.27$\pm$0.3 66.08$\pm$2.1 61.25$\pm$1.8 ()($\uparrow$)( ↑ ) 81.07$\pm$1.7 ()($\uparrow$)( ↑ ) 79.26$\pm$1.9 ()($\uparrow$)( ↑ ) 74.05$\pm$2.4 52.33$\pm$0.9 78.16$\pm$1.3
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{N},\cdot),\Phi_{\text{val}}(\mathbf{L},\cdot))$ 83.27$\pm$0.3 66.08$\pm$2.1 64.46$\pm$1.2 80.31$\pm$1.4 79.14$\pm$1.1 ()($\uparrow$)( ↑ ) 70.84$\pm$2.3 52.33$\pm$0.9 74.54$\pm$1.9 ()($\uparrow$)( ↑ )
$\mathtt{att}(\Phi_{\text{ind}}(\mathbf{L},\cdot),\Phi_{\text{val}}(\mathbf{N},\cdot))$ 82.53$\pm$1.9 ()($\uparrow$)( ↑ ) 65.73$\pm$2.4 62.31$\pm$1.4 79.54$\pm$2.1 ()($\uparrow$)( ↑ ) 80.51$\pm$1.2 74.05$\pm$2.4 51.37$\pm$0.5 ()($\uparrow$)( ↑ ) 78.16$\pm$1.3
Table 14: Graph classification performance using Pooling-NLSFs with uniform sub-bands. The symbol ()($\uparrow$)( ↑ ) denotes an improvement using Laplacian attention.
MUTAG PTC ENZYMES PROTEINS NCI1 IMDB-B IMDB-M COLLAB
$\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{L},\cdot)$ 86.41$\pm$1.9 68.76$\pm$0.9 69.88$\pm$1.3 84.27$\pm$1.3 80.33$\pm$0.8 60.40$\pm$1.3 51.01$\pm$1.3 71.28$\pm$0.9
$\Theta_{\text{ind}}^{\mathtt{P}}(\mathbf{N},\cdot)$ 84.52$\pm$0.8 67.19$\pm$1.2 66.37$\pm$2.1 81.12$\pm$1.7 77.05$\pm$2.2 62.08$\pm$1.6 52.48$\pm$1.0 72.03$\pm$1.2
$\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{L},\cdot)$ 83.49$\pm$1.1 67.12$\pm$1.0 65.38$\pm$1.7 82.89$\pm$2.2 80.57$\pm$2.4 72.85$\pm$1.4 50.91$\pm$1.4 75.02$\pm$1.2
$\Theta_{\text{val}}^{\mathtt{P}}(\mathbf{N},\cdot)$ 83.34$\pm$1.3 64.22$\pm$0.9 62.24$\pm$0.7 80.03$\pm$0.4 80.90$\pm$1.6 72.49$\pm$1.1 52.08$\pm$0.9 80.89$\pm$2.2
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{L},\cdot))$ 86.41$\pm$1.9 68.76$\pm$0.9 69.88$\pm$1.3 84.27$\pm$1.3 80.33$\pm$0.8 72.85$\pm$1.4 51.01$\pm$1.3 76.13$\pm$0.7 ()($\uparrow$)( ↑ )
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{N},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ 84.52$\pm$0.8 67.19$\pm$1.2 67.13$\pm$2.2 ()($\uparrow$)( ↑ ) 82.53$\pm$1.1 ()($\uparrow$)( ↑ ) 80.90$\pm$1.6 72.49$\pm$1.1 53.16$\pm$0.8 ()($\uparrow$)( ↑ ) 80.89$\pm$2.2
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{N},\cdot),\Theta_{\text{val}}(\mathbf{L},\cdot))$ 84.52$\pm$0.8 68.01$\pm$1.4 ()($\uparrow$)( ↑ ) 66.44$\pm$0.9 ()($\uparrow$)( ↑ ) 83.52$\pm$2.4 ()($\uparrow$)( ↑ ) 80.57$\pm$2.4 72.85$\pm$1.4 52.48$\pm$1.0 77.42$\pm$0.7 ()($\uparrow$)( ↑ )
$\mathtt{att}^{\mathtt{P}}(\Theta_{\text{ind}}(\mathbf{L},\cdot),\Theta_{\text{val}}(\mathbf{N},\cdot))$ 86.41$\pm$1.9 68.76$\pm$0.9 69.88$\pm$1.3 84.27$\pm$1.3 80.90$\pm$1.6 72.49$\pm$1.1 53.48$\pm$1.2 ()($\uparrow$)( ↑ ) 81.12$\pm$1.4 ()($\uparrow$)( ↑ )
Our primary objective in this work is to introduce new GNNs that are equivariant to functional symmetries, based on a novel spectral domain transferable between graphs using analysis and synthesis. Our NLSFs in Sec. 3.4 consider filters $g_{j}$ that are supported on the dyadic sub-bands $\left[\lambda_{N}r^{S-j+1},\lambda_{N}r^{S-j}\right]$. The closer to the low-frequency range, the denser the sub-bands become. We illustrate an example of filters $g_{j}$ supported on dyadic sub-bands with $r=\frac{1}{2}$ and $S=4$ in Fig. 3, showing that sub-bands become denser as they approach the low-frequency range. Our primary goal is to highlight the unique aspects of our method rather than sub-band separation, which, while crucial, is an important consideration across spectral GNNs. Therefore, we also present uniform sub-bands, where filters $g_{j}$ are supported on the uniform sub-bands $\left[(j-1)\frac{\lambda_{N}}{S},j\frac{\lambda_{N}}{S}\right]$. Note that the modifications required for our NLSFs are minimal, and most steps can be seamlessly applied with filters supported on the uniform sub-bands.
We evaluate our NLSFs with uniform sub-bands on graph classification tasks. We consider the graph benchmarks as in Sec. 5, including five bioinformatics datasets: MUTAG, PTC, NCI1, ENZYMES, and PROTEINS, and three social network datasets: IMDB-B, IMDB-M, and COLLAB. Note the sub-bands only affect our Value-NLSFs, where Index-NLSFs remain the same as in Sec. 3.4.
We report the graph classification accuracy of Graph-NLSFs and Pooling-NLSFs using uniform sub-bands in Tab. LABEL:tab:graph_classification_nlsfp_uniform_graph_nlsf and Tab. LABEL:tab:graph_classification_nlsfp_uniform_pooling_nlsf, respectively, where the $\Phi_{\text{ind}}(\mathbf{L},\cdot)$, $\Phi_{\text{ind}}(\mathbf{N},\cdot)$, $\Theta_{\text{ind}}(\mathbf{L},\cdot)$, and $\Theta_{\text{ind}}(\mathbf{N},\cdot)$ are the index-based NLSFs and therefore they are the same as in Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf and Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling. Interestingly, the Laplacian attention does not yield significant improvements for either Graph-level NLSFs or Pooling-NLSFs when considering uniform sub-bands. Moreover, we observe that the graph classification performance is generally worse than when using the dyadic sub-bands reported in Tab. LABEL:tab:graph_classification_nlsfp_ablation_graph_nlsf and Tab. LABEL:tab:graph_classification_nlsfp_ablation_pooling. We emphasize the importance of considering the spectral support of filters $g_{j}$. Empirically, we found that using the dyadic grid is more effective, which is why we focus on it and report those results in the main paper. However, exploring other sub-bands remains an important task for future work. For instance, we plan to investigate sub-bands based on $\{ar^{j}-b\}_{j=1}^{K}$ for other choices of $21<r<$ 2 and $b>0$. In the limit when $r\rightarrow 1$, $a\rightarrow\infty$ and $b\rightarrow\infty$ this “converges” to the uniform grid.